### v1

# LLM-Resistant Math Word Problem Generation via Adversarial Attacks

Abstract
Large language models (LLMs) have significantly transformed the educational landscape. As current plagiarism detection tools struggle to keep pace with LLMs’ rapid advancements, the educational community faces the challenge of assessing students’ true problem-solving abilities in the presence of LLMs. In this work, we explore a new paradigm for ensuring fair evaluation—generating adversarial examples which preserve the structure and difficulty of the original questions aimed for assessment, but are unsolvable by LLMs. Focusing on the domain of math word problems, we leverage abstract syntax trees to structurally generate adversarial examples that cause LLMs to produce incorrect answers by simply editing the numeric values in the problems. We conduct experiments on various open- and closed-source LLMs, quantitatively and qualitatively demonstrating that our method significantly degrades their math problem-solving ability. We identify shared vulnerabilities among LLMs and propose a cost-effective approach to attack high-cost models. Additionally, we conduct automatic analysis on math problems and investigate the cause of failure to guide future research on LLM’s mathematical capability.  [^1] Data and code are available: https://github.com/ruoyuxie/adversarial_mwps_generation
LLM-Resistant Math Word Problem Generation via Adversarial Attacks
Roy Xie      Chengxuan Huang      Junlin Wang      Bhuwan Dhingra Duke University {ruoyu.xie, jonathan.huang, junlin.wang2}@duke.edu {bdhingra}@cs.duke.edu
1 Introduction
Recent advances in large language models (LLMs) have revolutionized the world of education, primarily due to the great improvements in their natural language generation and problem-solving capabilities. This has transformed how students access information and complete assignments, raising significant concerns among educators in accurately evaluating students’ true problem-solving abilities with the presence of such powerful tools (OpenAI, 2023; Kung et al., 2023; Callanan et al., 2023). While efforts like plagiarism detection exist (Kirchenbauer et al., 2023; Mitchell et al., 2023), their effectiveness in identifying LLM-generated content is limited (Liang et al., 2023; Chaka, 2023), underscoring the need for more advanced anti-plagiarism methods to match LLM advancements.
At the same time, adversarial attacks on LLMs have gained more attention due to increased awareness of the potential risks associated with LLMs. Most work on adversarial attacks focuses on developing prompts to elicit specific outputs from LLMs (Wang et al., 2021; Zhu et al., 2023). Recent work suggests that even the most powerful aligned LLMs are vulnerable to such attacks (Zou et al., 2023; Carlini et al., 2023). Hence, we might expect that as LLMs become stronger, detecting their outputs will become more difficult, but adversarial examples may still persist Ilyas et al. (2019).
To this end, we introduce a new paradigm for generating homework assignments that LLMs cannot solve by utilizing adversarial attacks. We focus on math word problems (MWPs), which present a unique and challenging intersection of language and mathematics. In this work, we ask the question: “Can LLMs still solve an MWP after changing its numeric values?” We aim to create MWPs that LLMs are unable to solve while maintaining the coherence and original difficulty of the problems. This goal diverges from typical adversarial attacks in machine learning, where small, often imperceptible changes are made to input data (images, text, etc.) to deceive models into making incorrect predictions. Here we do not focus on making imperceptible changes, but rather on preserving the coherence, style, and, most importantly, difficulty of the original problems. Our aim is not just to challenge LLMs but to do so in a way that reflects real-world educational standards to ensure the MWPs remain educationally valuable and relevant.
We generate adversarial examples by editing the numeric values in MWPs. While conceptually simple, doing this effectively is challenging for a couple of reasons: (i) the answer of the resulting problem changes and must be recomputed to assess the robustness of LLMs; and (ii) preserving the difficulty and validity of the modified problem requires us to ensure that all the intermediate calculations in the new problem are consistent with the original problem. For example, if the original problem did not have fractions as intermediate values, in an educational context we would not want the modified problem to have such values either. To tackle these challenges, we first convert MWPs to a code representation and leverage abstract syntax trees (ASTs) to systematically generate adversarial problems.
We evaluate several major LLMs on their math-solving ability under adversarial settings, demonstrating the effectiveness of our method with a significant degradation across various LLMs. We investigate universal attacks and attack transferability, proposing a cost-effective approach to attack high-cost models, achieving a 90% request reduction rate without sacrificing performance. We verify the validity of our generated problems by conducting human evaluations. Lastly, we conduct in-depth automatic analyses to investigate the mathematical vulnerabilities in LLMs and find, unsurprisingly, that the values of the answer and the number of operations have strong correlations with correctness.
Figure 1: Method Overview: given a MWP that an LLM can correctly solve, our method first transforms it into Python code. Next, the Python code is converted into an AST representation which is used to generate adversarial problems by modifying the numeric values in the problem. We place constraints on the nodes of the AST representation to ensure that the modified problem maintains the same difficulty level as the original problem. Despite this, we find that the resulting adversarial examples cause LLMs to predict incorrect answers.
2 Background and Related Work
Fair Evaluation for Educational Purpose
The increasing sophistication of LLMs has made it challenging for plagiarism detection tools to accurately identify content generated by machines (Chaka, 2023; Liang et al., 2023). Previous work suggests that detecting LLM-generated content may become nearly impossible as LLMs improve (Sadasivan et al., 2023). The issue extends beyond text-based content; various other disciplines are also threatened by the advancing capabilities of LLMs. For instance, LLMs can now provide step-by-step solutions for MWPs, which complicates the matter of fair evaluation, as students may pass off machine-generated solutions as their own without understanding the underlying mathematical concepts. It remains an open question of how to ensure fair evaluation for students in the presence of LLMs.
LLMs Math-solving Ability
The capability of LLMs to solve math problems is currently a hot topic (OpenAI, 2023; Yu et al., 2023; Xu et al., 2023). Studies found that LLMs can significantly improve their math-solving ability through prompt engineering (Yu et al., 2023; Imani et al., 2023), such as chain-of-thought (CoT) (Wei et al., 2022). Converting the MWPs’ solving steps into code or symbolic representations can also improves LLM’s performance (Li et al., 2023; He-Yueya et al., 2023; Gao et al., 2023). While different prompting strategies exist, our focus is to generate adversarial examples that LLMs are unable to solve. Therefore, we aim to minimize prompt engineering by using a unified prompt with zero-shot CoT for all queries in this work. [^2] Prompt template is in Appendix A.2.
Adversarial Attacks on MWPs
Adversarial attacks on LLMs focus on modifying prompts to cause LLMs to output desired content (Zhang et al., 2020; Zou et al., 2023; Carlini et al., 2023). Similarly, our focus is to modify the MWPs to cause LLMs to output incorrect answers. Bubeck et al. (2023) conducts a limited number of memorization tests on MWPs by randomly changing the numeric values between -10 and 10, suggesting that state-of-the-art LLMs do not simply rely on memorizing problems but on applying a general solution method. However, the number of examples in Bubeck et al. (2023) is relatively small to reach such a conclusion, and we will show in our work that modifying numbers in MWPs does cause LLMs to fail on a larger scale. Zhou et al. (2023), on the other hand, rephrases MWPs by changing the words within the problem and keeping all the numeric values the same. However, this approach makes it challenging to preserve the problem’s original context and structure, and sometimes could introduce inconsistent and problematic content, as shown in Appendix B. Additionally, Zhou et al. (2023) also requires extensive human validation to ensure modification quality. Therefore, we focus on changing numerical values, which ensures the preservation of the underlying logic. To the best of our knowledge, this is the first study to undertake such an endeavor.
3 Methodology
MWPs are presented using natural language, which creates challenges for systematic structural and syntactic modifications. In this section, we describe our approach to generating adversarial MWPs. An overview of our method can be found in Figure 1.
3.1 Preliminaries
We denote an original problem-answer pair as $p=(x,y)$, where $x$ is a sequence of tokens $x_{1},\dots,x_{n}$ in the problem and $y$ is its ground-truth answer. We denote the set of all possible adversarial modifications to it as:
$A(x,y)=\{(\tilde{x},\tilde{y}):\tilde{x}_{i}\in F(x),\tilde{y}=G(\tilde{x})\}.$ (1)
Here $G$ is a ground-truth function that computes the correct answer for a math problem (in our case this is generated Python code). $F$ is a function which maps the elements of a sequence as: $F(\tilde{x}_{i})\sim\mathbb{R}$ (i.e., a random real number) if $x_{i}$ is numeric and $F(\tilde{x_{i}})=x_{i}$ otherwise. This entire set of adversarial examples will also consist of many unnatural and difficult problems, hence later in §3.3 we will introduce filtering constraints for selecting the adversarial examples that preserve difficulty.
3.2 Mapping From MWPs to Code to Tree
To structurally modify MWPs, we first convert MWPs into Python code by prompting GPT-4-turbo to generate the code based on the problem-solving steps provided by the datasets. [^3] Prompt can be found in Appendix A.1. As a sanity check, we first select 200 MWPs each from GSM8K (Cobbe et al., 2021) and MultiArith (Roy and Roth, 2015) and check if the generated program results in correct answers. Given the execution steps in code are almost always deterministic, encouragingly, we found only 2 and 5 generated codes have incorrect logic or answers in GSM8K and MultiArith, respectively. We further investigate the generated Python code in §5.3.
Next, we utilize AST, a tree representation for the abstract syntactic structure of programming language code, to convert the generated Python code into a tree. We build the ASTs by traversing through the Python code and constructing nodes corresponding to each statement. The final print statement that outputs the answer is the root of the tree. There are mainly two types of nodes: operation nodes, which carry out the operations in the tree and are the non-leaf nodes, and variable nodes, which correspond to the numeric values from the problem and are the leaf nodes. We discuss the nodes in detail in Appendix C.
3.3 Adversarial Example Generation
In an adversarial setting, attackers can perform multiple attacks on the system. Similarly, we generate multiple adversarial examples which are different versions of the original MWPs to attack LLMs.
Educational Context
It’s essential to maintain the original difficulty and coherence of the MWPs despite changes in their numeric values. For instance, changing one-digit multiplication to four-digit multiplication significantly alters the problem’s complexity. Furthermore, to keep the mathematical logic intact, the order of magnitude of numbers in the problem should remain unchanged. For example, altering a problem from “subtracting 2 apples from 5” to “subtracting 4 apples from 2” introduces the concept of negative numbers, potentially misaligning with the original problem’s context. Lastly, changes should be contextually coherent. For instance, a generated problem with “the first half of 3 people” would be confusing.
Filtering Constraints
To minimize the difference between the original MWP and its adversarial counterpart while adhering to the educational context, we define a set of Boolean constraints for each node to control the generation quality. Given the original node value $h$ and the new value of node $h^{\prime}$, a newly generated problem is valid if and only if the $h^{\prime}$ satisfies the same set of constraints that $h$ has, for all nodes. A list of node constraints is:
  • Positivity: if $h$ is positive, then $h^{\prime}$ should be positive. This constraint avoids generating phrases like “get 5 apples from 2” since this would produce a negative node.
  • Integer: if $h$ is an integer, then $h^{\prime}$ should remain an integer. This ensures that phrases like “the first half of 3 people” wouldn’t appear, since the original problem likely has an integer value in the intermediate node.
  • Proper Fraction: if $h$ is between 0 and 1, then $h^{\prime}$ should be between 0 and 1. This prevents phrases such as “John eats 4/3 of his chimichanga” and “the bucket is filled till 150% full” from being produced since the generated problems wouldn’t make logical sense for many problems when a number is no longer a proper fraction.
Constrictive Generation Methods
The node constraints only ensure that the adversarial examples are valid in the numerical sense; instances that don’t make logical sense like “$28$-hour work day”, which is not common, can still be generated. Furthermore, different values for the variable nodes can lead to vastly different difficulty levels, as $h^{\prime}$ can be significantly larger than $h$.
To mitigate this, we propose three generation methods to distinguish generated adversarial examples with varying levels of difficulty. We decompose the values $h\approx a\times 10^{b}$, where $a$ is an integer that is not divisible by 10 and with at most $c$ digits. Depending to the numbers $h$, $a$, and $b$, the generation methods are defined as follows:
  • $M1$: Free Generation: regardless of the value of $a$, $a^{\prime}$ is drawn from a uniform distribution between $1$ and $10^{c}$. If the variable node $v$ is a divisor of a division node, then $a^{\prime}$ is drawn from a uniform distribution between $1$ and $\left\lfloor 10^{c/2}\right\rfloor$. The newly generated value $h^{\prime}$ will be $a^{\prime}\times 10^{b}$. The reason for picking smaller numbers for a divisor is that the probability of picking a number from $1$ and $10^{c}$ that is divisible for another number from $1$ and $10^{c}$ is very low for big values of $c$.
  • $M2$: Count of Digits: this generation method ensures that the newly generated value $h^{\prime}$ has the same number of digits as $h$. For example, for $h=100$, some possible $h^{\prime}$ are 942, 589, or 264. This ensures $h^{\prime}$ is always within a reasonable range relative to $h$ and maintains similar problem difficulty levels. For more variability, $h^{\prime}$ can range from 1 to 99 if $h$ has only one digit. For decimal variable values like $h=1.25$, we will use $a=125$ to generate numbers like $a^{\prime}=473$, then convert them back to $h^{\prime}=4.73$.
  • $M3$: Count of Scientific Numbers: this generation method tries to ensure that the new value shares a similar scientific digit count and is within a similar range to the original value. For example, for $h=1500$, $h^{\prime}$ can be 1700, 800, 1200. The rationale is that changing a number to a larger one does not necessarily make calculations harder, but more scientific numbers do. An example would be comparing the numbers 150,000 and 172,568.
$M3$ is the most restrictive generation, followed by $M2$, and $M1$ is the least restrictive, therefore:
$M1(A(x,y))\supseteq M2(A(x,y))\supseteq M3(A(x,y)).$ (2)
The more restrictive a method is, the closer the difficulty levels and coherence are between the adversarial examples and the original problem. Therefore, we use $M3$ as our main generation method as it ensures the adversarial examples comply with the constraints while also gauging the difficulty of problems. Although generations from $M1$ and $M2$ may not fit in the educational context, we are interested in studying LLMs’ math-solving abilities by simply changing the numeric values. We discuss how different methods impact model performance in §4.4. We describe each generation method in more detail in Appendix D and present some sample generations based on each method in Table 5.
4 Experiments and Results
In this section, we present multiple experiments and demonstrate the effectiveness of our method on attacking various LLMs.
4.1 Experimental Setup
Datasets
We generate adversarial examples from GSM8K and MultiArith. Both datasets are commonly used for evaluating LLMs’ mathematical capabilities, and MultiArith is sourced from math worksheets used by elementary school students to practice math problems (Roy et al., 2015).
Models
We conduct experiments on 7 open-source models: MetaMath 7B, 70B (Yu et al., 2023), Mistral 7B (Jiang et al., 2023), Llama-2 13B (Touvron et al., 2023), WizardMath 13B (Xu et al., 2023), Vicuna 13B (Chiang et al., 2023), CodeLlama (Roziere et al., 2023), and 2 closed-source models: GPT-4-Turbo (OpenAI, 2023) and GPT-3.5-Turbo (OpenAI, 2022). [^4] Specifically, we use GPT-4-0125-preview and GPT-3.5-turbo-0125. For simplicity, we refer to them as GPT-4 and GPT-3.5 for the rest of the paper. The model selection was largely based on the LLMs’ math-solving ability. Our intention is to evaluate a wide range of LLM performances, including the math-tuned LLMs such as MetaMath and WizardMath, popular open-source LLMs such as Llama-2 and Mistral, and the state-of-the-art API-based GPT models.
Metrics
Given a set of original problem-answer pairs $P$ and a LLM $L$, we use the follwing evaluation metrics:
  • Original Accuracy (OA): the accuracy of $L$ on the original problems:
    $\text{OA}(L)=\frac{\sum_{(x,y)\in P}\mathbf{1}\{L(x)=y\}}{|P|}.$ (3)
  • Attack Accuracy (AA): given an indicator function $I_{xy}$:
    $I_{xy}=\mathrm{1}\left[\forall(\tilde{x},\tilde{y})\in A(x,y):L(\tilde{x})=\tilde{y}\right],$ (4)
    $L$’s accuracy on the adversarial examples:
    $\text{AA}(L)=\frac{\sum_{(x,y)\in P}I_{xy}}{|P|}.$ (5)
  • Attack Success Rate (ASR): the relative decrease in accuracy due to adversarial modifications:
    $\text{ASR}(L)=\frac{\text{OA}(L)-\text{AA}(L)}{\text{OA}(L)}.$ (6)
4.2 Rephrasing Attacks
We examine the effectiveness of a commonly used rephrasing approach, where the modifications are centered around the words within the MWPs. Zhou et al. (2023) freeze the logical entities (e.g., entity names and values) and iteratively go through each unfrozen token to swap it with similar tokens that could cause LLMs to fail. They then validate the adversarial example with human manual checking to ensure the example’s quality. Zhou et al. (2023) present the RobustMath dataset, which includes 214 original problems from a combination of GSM8K and MultiArith, and generate 300 rephrased versions. As a baseline, we report their ASR in Table 2. We observe that rephrasing attacks result in lower accuracy for 4 out of the 7 models; however, some models show improved accuracy, suggesting that such attacks may not generalize across different LLMs.
MultiArith GSM8K
$M3$ $M2$ $M1$ $M3$ $M2$ $M1$
Model OA AA ASR AA ASR AA ASR OA AA ASR AA ASR AA ASR
Mistral 7B 37 0 100 0 100 0 100 29 0 100 0 100 0 100
MetaMath 7B 100 74 26 10 90 0 100 95 28 71 9 91 0 100
Llama 2 13B 12 0 100 0 100 0 100 10 0 100 0 100 0 100
WizardMath 13B 89 20 77 5 94 0 100 89 11 88 2 98 0 100
Vicuna 13B 76 4 95 1 99 0 100 60 0 100 0 100 0 100
CodeLlama 34B 11 0 100 0 100 0 100 6 0 100 0 100 0 100
MetaMath 70B 99 86 13 30 70 0 100 98 50 49 17 83 0 100
GPT-3.5-turbo 97 74 24 47 52 0 100 91 52 43 31 66 0 100
Table 1: Main Results: we generate adversarial examples based on the given constraint. Overall, simply changing the numeric values in the original math problems causes a consistent performance drop across all models, even with the most restrictive generation method.
4.3 Number of Adversarial Examples
We investigate how the number of adversarial examples impacts model performance. Figure 2 shows the average number of possible adversarial examples per question for each constraint.
Figure 2: Possible Generations: with 30,000 attempts for each given constraint, we calculate the average number of adversarial examples generated for each problem.
Figure 3: Increase Examples: we report the model performance given various number of adversarial examples. For several models, only 10 adversarial examples are enough to degrade the performance.
While on average each problem generates over a thousand modifications, it is worth noting that the actual number of possible generations varies significantly based on the values and the number of variable nodes in a problem. For example, a problem with only one digit value and one variable node has a potential maximum of 9 possible modifications under the $M3$ generation method. Given this reason and computational cost considerations, in this work we randomly select 100 problems from MultiArith and GSM8K separately and generated 100 adversarial examples for each problem. Despite the seemingly small number of problems, the creation of 30,000 adversarial examples per dataset given three different generation methods (totaling 60,000 examples across both datasets) is significant. We present the GSM8K results in Figure 3, observing a clear pattern of accuracy decline across all models as the number of examples increases. Interestingly, for several models, just 10 adversarial examples are enough to degrade the performance.
4.4 Main Results
We evaluate each adversarial example for each problem against all models given different generation methods. A problem is considered incorrect if there exists at least one adversarial example under that problem that results in an incorrect answer. We refer such problem as successful attack. Here we do not run the full generated adversarial examples against GPT-4 due to cost consideration. Instead, we propose a cost-effective way to query expensive models in §4.6. We report the model performance with all three generation methods in Table 1. Ranging from the most lenient method, $M1$, where the numbers could be fairly wild, to the most stringent condition, $M3$, we observe a consistent accuracy drop across all models, even for the strong models (e.g., math-tuned, larger size, or API-based). For the most original-problem-like generation, $M3$, we observe that the weaker models (e.g., smaller size or general-purpose) fail to generate any correct answers in every case. As a comparison to rephrasing attack, we present the average ASR of $M3$ in Table 2. Even as the most restrictive generation method, $M3$ significantly outperforms RobustMath in every model.
Model ASR (RobustMath) ASR (ours)
Mistral 7B 0.0 100.0
MetaMath 7B 13.0 48.5
Llama-2 13B 0.0 100.0
WizardMath 13B 1.0 82.5
Vicuna 13B 0.0 97.5
CodeLlama 34B 67.1 100.0
MetaMath 70B 11.1 31.0
GPT-3.5-Turbo 16.9 33.5
Table 2: Comparison Result: given that RobustMath comprises MultiArith and GSM8K, we calculat the average ASR of $M3$ from our results for MultiArith and GSM8K. Our method significantly outperforms RobustMath in every model.
4.5 Universal Attacks
We are interested in whether there exist problems that all LLMs fail, also known as universal attacks. We count the number of successful attacks from each model and calculated the percentage of common successful attacks among them. We present the results in Table 3 and observe a clear pattern of decreasing universal attacks with an increased number of models. This suggests that diversity among models can lead to increased resilience against such attacks. We also observe that the impact of universal attacks varies significantly across different generation methods, with $M3$ having the lowest number of universal attacks.
Model Ct. M3 (%) M2 (%) M1 (%)
Mistral 7B 1 100.0 100.0 100.0
MetaMath 7B 2 70.0 90.0 100.0
Llama-2 13B 3 67.0 90.0 100.0
WizardMath 13B 4 52.0 80.0 100.0
Vicuna 13B 5 44.0 78.0 100.0
CodeLlama 34B 6 21.0 67.0 99.0
MetaMath 70B 7 9.0 49.0 87.0
GPT-3.5-Turbo 8 9.0 46.0 83.0
Table 3: Universal Attack: the number of universal attacks decreases as the number of models increases, implying that diversity among models enhances resilience against such attacks.
4.6 Efficient Attacks
Figure 3 shows that scaling up the number of adversarial examples results in a steady performance drop for all models. However, this scaling might not be viable for API-based models such as GPT-4 or GPT-3.5 due to request limits and high costs. To circumvent this, we propose an efficient attack method for API-based models by using adversarial examples in a targeted manner. The idea is simple: we attack model $A$ (target model, GPT-4 in our case), with successful attacks from a cheaper model $B$ (e.g., open-source, lower API cost). We first select 50 problems along with their 100 adversarial examples running against GPT-4. Next, we select 50 successful attacks from a cheaper model and run them against GPT-4. Note that for some models, there are no correct responses, even with $M3$. Therefore, we only compare $M3$ with the models that has correct answers. We compare the results in Table 4 and observe a significant request reduction while achieving similar ASR.
Method Req. RR. (%) ASR (%)
Original 5000 0 10
MetaMath 7B 1389 72 8
WizardMath 13B 1745 65 6
MetaMath 70B 672 87 8
GPT-3.5-turbo 456 91 10
Table 4: Efficient Attack: a targeted approach to attack high-cost models. We leverage successful attacks from cheaper models to attack GPT-4, achieving the same ASR with about 90% reduction in requests.
4.7 Human Evaluation
We conduct a human evaluation on the generated adversarial examples to ensure they are indeed valid. We randomly select one adversarial example from 20 problems from all three generation methods (totaling 60 examples) that GPT-3.5 failed on for GSM8K and presented them to three human evaluators to assess (i) correctness: whether the answer to the modified problem is correct; (ii) coherence: whether the problem is contextually coherent; and (iii) similarity: whether the newly generated problem’s difficulty level is similar to the original one. The annotators were asked to make a binary decision to determine whether the criteria described above were met. We present the average of their scores in Figure 4. Overall, we found that all human evaluators agreed with the correctness. $M3$ scores highly across all metrics, indicating that it indeed preserves the original problem’s coherence and difficulty level. On the other hand, the scores for coherence and similarity varies between generation methods, suggesting that while they all successfully perturbed the model, some of the examples may not be useful in an educational context.
Figure 4: Human Evaluation: we report the average score among three annotators on the successful attacks from GPT-3.5-turbo. $M3$ shows the highest score across all metrics, indicating our best generation method correctly generate contextual coherent problems that preserve the original level of difficulty.
5 Analysis and Discussion
In this section we conduct several analyses to better understand our generation methods and LLMs mathematical capabilities.
5.1 Transferability
We investigate the transferability of adversarial examples among the models. We attack model $A$ with the successful attacks from model $B$ and calculate the number of common successful attacks between these two models. We present the GSM8K $M3$ result in Figure 5. Interestingly, we observe that Vicuna 13B, Llama2 13B, CodeLlama 34B, and Mistral 7B exhibit a high percentage of transferability, reaching 100%, indicating significant vulnerability and a strong correlation among these models. On the other hand, comparisons with MetaMath 70B and GPT-3.5 tend to have a lower transferability rate. While it might suggest a form of resistance to adversarial examples that affect other models, it is not a measure of robustness as it could be that those models fail on entirely different examples. This finding suggests that certain LLMs might struggle with particular numbers or patterns. To further understand these phenomena, we conduct further analysis in §5.2.
Figure 5: Transferability: we present the adversarial example transferability amount all models by comparing each model against all other models. Compare to the math-tuned and production models, the weaker models such as Mistral 7B and LLama2 13B exhibit significant vulnerability and a strong correlation amount them.
5.2 Question Feature Analysis
As a step towards understanding problems that are hard for LLMs, we conduct regression analysis to understand the effects of various features of the problems. More specifically, we construct features from the 20,000 MWPs using the generation method $M3$ (200 problems, each with 100 modifications). The input features include operation counts, value ranges, and node counts in the AST, resulting in a total of 51 input features. The features are plotted against the correctness of the model, with correct being $1$ and wrong being 00.
We present the coefficients of the eight models with some selected features in Table 7 in the Appendix. The coefficients are only compared within each model, but not across models. A more suited metric to compare between models is accuracy, such as Table 1 and Figure 3. We present a set of graphs showing models’ accuracy versus some prominent features in Figure 7. Overall, we found that the value of the answer has a stronger correlation to the correctness. Somewhat surprisingly, most models have positive correlations between the number of constants and variables in a question and the correctness. The reason might be that the problems with more numbers have more straightforward operations (additions and subtractions). Additionally, we found that most high-performing models exhibit a negative correlation with the number of divisions in the problem, except for MetaMath 70B, and operation counts negatively correlate with correctness for all models.
5.3 Code Analysis
While the majority of generated Python code adheres to correct problem-solving logic and produces answers correctly, to ensure that the generated code is correctly being converted into ASTs, we manually validate the generated ASTs and identify the following errors: (1) Incorrect answers or code; (2) Use of loops, conditional, and comparison statements, and functions; (3) Use of expressions such as x = y // z + (y % z > 0); (4) Number misalignment, where the same number appears multiple times in the code or the problem. For GSM8K, 7 out of 200 problems were filtered out. For MultiArith, 15 out of 200 problems were filtered out. We discuss the methods of identifying and handling such errors in Appendix F.
6 Conclusion
In conclusion, we introduce a novel approach to generate MWPs that challenge the LLM’s math-solving abilities, thereby addressing the urgent need for methods that ensure fair educational evaluations. By utilizing AST to create adversarial examples, our generation method significantly degrades both open- and closed-source LLMs without compromising the problems’ original difficulty and coherence. We investigate the shared mathematical vulnerabilities among LLMs and propose a cost-effective method to attack high-cost API-based models, achieving up to a 90% decrease in request rates without compromising performance. Our automatic analyses shed light on the specific mathematical aspects that most influence LLM performance. This work not only paves the way for further research into developing more robust and ethically sound educational tools but also offers valuable insights into LLMs’ limitations, contributing to the ongoing discourse on the ethical use of LLMs in education.
7 Limitations
Perceived Difficulty Level: In this work, the correlation between the complexity of problems generated to challenge LLMs and the actual difficulty perceived by human students has not been empirically validated. It is plausible that modifications considered neutral or simple for LLMs could be perceived as more challenging by human students. We plan to further investigate this in future work.
Imperfect Logical Coherence: While $M3$ produces similar numbers to the original problems, logical coherence is not perfect in few cases. For example, the phrases like “25 hour in a day” and “37 days in a month”. While this issue can be easily addressed by employing heuristic rules to associate certain numbers (e.g., 7, 24, 365) with specific entities (e.g., “days”, “hours”, “year”), we try to keep everything as generalized as possible without overfitting to specific cases.
Cannot Alter Words: In this work, we only consider actual numeric values in a question; however, there are problems where words also represent numeric values. For example, a question from GSM8K is “A frog lays her eggs over a series of 4 days. The first day she lays 50 eggs. The second day, she doubles her production of eggs. The third day she lays 20 more than the second day, and the last day she doubles the first three days total. How many eggs did the frog lay over the span of the 4 days?” contains words like “doubles” and “three” that also represent numeric values. In the future, we plan to add a preprocessing step for the problems to extract as many numeric values from a question as possible, resulting in more diverse generations of problems.
8 Ethics Statement
While our primary intention is to address concerns about academic dishonesty, we acknowledge that our strategy might also lead to potential exacerbating educational inequality. By designing math problems specifically to be unsolvable by LLMs, the educators without access to resources, training, or the necessary tools could be placed at a disadvantage. This approach risks amplifying the disparities between institutions or individuals with differing levels of technological access.
9 Acknowledgements
We are very grateful to Sandy Yeung and Jack Goffinet for the helpful initial project discussion. Thank you to the anonymous reviewers for their feedback. The research reported here was supported by the Learning Engineering Virtual Institute funded by leading education philanthropists and organizations through Grant G-23-2137070, to the University of Florida and its partner institutions. The opinions expressed are those of the authors and do not represent the views of the University of Florida, the partner institutions, or those of the philanthropists and organizations.
Appendix A Prompt Template
A.1 Code Generation Prompt
[…]
A.2 Zero-shot CoT
[…]
Appendix B Rephrasing Example
Figure 6: A rephrasing attack example from Zhou et al. (2023). The left side shows the original problem, and the right side is the rephrased version. Although planes and helicopters are conceptually similar, the maximum flying distance for a helicopter is usually between 300 to 400 miles. The rephrasing introduces a subtle and incorrect factual error into the problem, which could be hard to detect at times.
Appendix C AST Nodes
With the two main types of nodes, we further distinguish them into following nodes:
  • Binary operation node: A node consists of two operands (nodes) and one main operation.
  • Unary operation node: A node consists of only one operand and one main operation.
  • Variable node: Represents the corresponding variable and its value from the problem. The set of variable nodes will be represented by $V=\{v_{1},\dots,v_{n}\}\subset S$, where $S$ represents the full tree.
  • Constant node: A special type of variable node, where the number appearing in the code does not correspond to any number in the problem. For example, a problem mentioning “one year” implies a number of “365” days in the code.
We associate the number in the original question with the variable node so that a change in the variable node will also change the number in the question. If a number appears multiple times in the code and the problem, then we associate them based on the specific procedures described in Section F.4.
Appendix D Generation Details
In this section, we describe the generation detail for new numbers in each generation method. Each variable node in an AST generates a new number $val(v^{\prime})=p^{\prime}$ by following Algorithm 2, which takes the AST $S$, a maximum number of attempts $N$, a maximum number of scientific digits $c$, the set of Boolean constraints $C$, and the desired generation $M$ method as inputs. It produces $m$ sets of new problems $W^{\prime}=\{V^{\prime}_{1},\dots,V^{\prime}_{m}\}$ that satisfy the given node constraints defined in §3.3.A newly generated adversarial example is valid if all nodes have their given constraints satisfied. Unless otherwise stated, we use $N$=30,000 and $c$=6 for all experiments.
Input: $v_{i},M,c$
Output: $v^{\prime}_{i}$
[…]
      if $v_{i}$ is a divisor for a division node then
           $a^{\prime}_{i}\sim\text{Uniform}(1,10^{c/2})$
           else
                $a^{\prime}_{i}\sim\text{Uniform}(1,10^{c})$
[…]
                     Let $d$ be the number of digits $val(v_{i})$ has. if $d=1$ then
                          $a^{\prime}_{i}\sim\text{Uniform}(1,99)$
                          else
                               $a^{\prime}_{i}\sim\text{Uniform}(10^{d-1},10^{d}-1)$
                               $b\leftarrow 0$ if $b>0$ else $b$
[…]
                                    if $1\leq a_{i}\leq 9$ then
                                         $a^{\prime}_{i}\sim\text{Uniform}(1,9)$
                                         else
                                              $a^{\prime}_{i}\sim\text{Pois}(a_{i})$
$val(v^{\prime}_{i})\leftarrow a^{\prime}_{i}\times 10^{b}$ return $v^{\prime}_{i}$
Algorithm 1 Number Generation Based on Method
Input: $S=\{s_{1},\dots,s_{u}\}$, $V=\{v_{1},\dots,v_{n}\}\subset S$, $N$, $c$, $C$, $M$
Output: $W^{\prime}=\{V^{\prime}_{1},\dots,V^{\prime}_{m}\}$, where $V^{\prime}_{j}=\{v^{\prime}_{1},\dots,v^{\prime}_{n}\}$
$attempts\leftarrow 0,j\leftarrow 1$ $W^{\prime}=\emptyset$ while $attempts<N$ do
      $V^{\prime}_{j}=\emptyset$ for $v_{i}\in V$ do
           Obtain $v^{\prime}_{i}$ from Algorithm 1 with input $v_{i}$, $M$, and $c$ $V^{\prime}_{j}\leftarrow V^{\prime}_{j}\cup\{v^{\prime}_{i}\}$
           $attempts\leftarrow attempts+1$ $S^{\prime}\leftarrow S$ Apply the new sequence of values from the variable nodes $V^{\prime}_{j}$ to $S^{\prime}$ $accept\leftarrow True$ for each node $s^{\prime}_{k}$ in $S^{\prime}$ do
[…]
                     $accept\leftarrow False$
                     if accept then
                          $W^{\prime}\leftarrow W^{\prime}\cup\{V^{\prime}_{j}\}$ $j\leftarrow j+1$
return $W^{\prime}$
Algorithm 2 Problems Generation
Method Question Answer
Original
Sarah makes 5 times more money per hour than
Connor does. If Connor earns \$7.2 per hour, how
much does Sarah make in an 8-hour day?
288
M3
Sarah makes 7 times more money per hour than
Connor does. If Connor earns \$8.0 per hour, how
much does Sarah make in an 7-hour day?
392
M2
Sarah makes 20 times more money per hour than
Connor does. If Connor earns \$7.7 per hour, how
much does Sarah make in an 8-hour day?
1232
M1
Sarah makes 76780 times more money per hour than
Connor does. If Connor earns \$7655.5 per hour, how
much does Sarah make in an 12568-hour day?
7387335796720
Original
Mary does her grocery shopping on Saturday. She
does her shopping only at a specific store where
she is allowed a credit of \$100, which must be paid
in full before her next shopping trip. That week she
spent the full credit limit and paid \$15 of it on
Tuesday and \$23 of it on Thursday. How much
credit will Mary need to pay before her next
shopping trip?
62
M3
Mary does her grocery shopping on Saturday. She
does her shopping only at a specific store where she
is allowed a credit of \$80, which must be paid in full
before her next shopping trip. That week she spent
the full credit limit and paid \$12 of it on Tuesday and
\$19 of it on Thursday. How much credit will Mary
need to pay before her next shopping trip?
49
M2
Mary does her grocery shopping on Saturday. She
does her shopping only at a specific store where she
is allowed a credit of \$432, which must be paid in
full before her next shopping trip. That week she
spent the full credit limit and paid \$91 of it on
Tuesday and \$76 of it on Thursday. How much
credit will Mary need to pay before her next
shopping trip?
265
M1
Mary does her grocery shopping on Saturday. She
does her shopping only at a specific store where she
is allowed a credit of \$56347, which must be paid in
full before her next shopping trip. That week she
spent the full credit limit and paid \$54731 of it on
Tuesday and \$1566 of it on Thursday. How much
credit will Mary need to pay before her next
shopping trip?
50
Original
A bird watcher records the number of birds he sees each
day. One Monday he sees 70 birds. On Tuesday he sees
half as many birds as he did on Monday. On Wednesday
he sees 8 more birds than he did on Tuesday. How many
total birds did the bird watcher see from Monday to
Wednesday?
148
M3
A bird watcher records the number of birds he sees each
day. One Monday he sees 80 birds. On Tuesday he sees
half as many birds as he did on Monday. On Wednesday
he sees 2 more birds than he did on Tuesday. How many
total birds did the bird watcher see from Monday to
Wednesday?
162
M2
A bird watcher records the number of birds he sees each
day. One Monday he sees 26 birds. On Tuesday he sees
half as many birds as he did on Monday. On Wednesday
he sees 39 more birds than he did on Tuesday. How many
total birds did the bird watcher see from Monday to
Wednesday?
91
M1
A bird watcher records the number of birds he sees each
day. One Monday he sees 57010 birds. On Tuesday he sees
half as many birds as he did on Monday. On Wednesday he
sees 86391 more birds than he did on Tuesday. How many
total birds did the bird watcher see from Monday to
Wednesday?
200411
Table 5: Three examples of questions generated by the three generation methods.
Appendix E Rephrasing Attack Performance
Model OA AA ASR
Mistral 7B 10.3 18.7 0.0
MetaMath 7B 91.1 79.3 13.0
Llama-2 13B 2.3 8.3 0.0
WizardMath 13B 71.0 70.3 1.0
Vicuna 13B 46.3 51.7 0.0
CodeLlama 34B 31.3 10.3 67.1
MetaMath 70B 93.0 82.7 11.1
GPT-3.5-Turbo 91.1 75.7 16.9
Table 6: Rephrasing Attack: As a baseline, we present the performance of RobustMath Zhou et al. (2023). While rephrasing the problem results in lower accuracy for 4 out of 7 models, a few models show improved accuracy with the rephrased version. This suggests that rephrasing the MWPs may not generalize effectively across different LLMs.
Appendix F Code Analysis
F.1 Incorrect Answer from the Dataset
There are instances where the dataset-provided answers might be wrong. An example of such an instance is the following question from the MultiArith dataset: "Paige’s team won their dodgeball game and scored 41 points total. If Paige scored 11 of the points and everyone else scored 6 points each, how many players were on her team?", which the dataset provided 5 as the answer. However, the correct answer should be 6, as Paige is on the team as well. The code provided by GPT-4 counted Paige to the team. To automatically detect such instances, the value of the final answer node will be compared to the answer given by the dataset, and the question will be filtered out for further generations of problems if the two answers don’t match.
F.2 Incorrect Answer from GPT-4
Similarly, there are also instances where GPT-4 generated codes that provided the wrong answer to the given question. One such question is from the GSM8K dataset: "Jen decides to travel to 3 different countries. He has to pay \$400 for the supplies he needs, in total. The tickets for travel cost, in total, 50% more than the supplies. How much does travel cost?" The correct answer should be $400+400\times 1.5=1000$, but the code generated by GPT-4 outputted \$600. Similar to the previous error, instances where the code outputs the wrong answers will be detected by comparing them to answers given by the dataset, and the instances will be discarded from generating math problems.
F.3 Contains Control Flow Statements
Although specified in the prompt, GPT-4 might generate codes that include control flow statements (including "if", "for", and "while" statements). Codes that contain control flow statements are considered errors and discarded because the codes are usually only applicable to the original instance of the question. Furthermore, the code might not halt for some combinations of numbers, resulting in no answers associated with the generated instances.
F.4 Number Misalignment
To ensure the correct generation of new questions, numbers in the original question need to be associated with numbers in the code. Given the sequence of numbers in the question $(q_{1},\dots,q_{n})$ and the sequence of numbers in the code $(c_{1},\dots,c_{m})$, the goal of grounding is to associate $q_{i}$ to the correct $c_{j}$ in the code. For distinct $q_{i}$’s, such pairing is trivial. However, there are instances where some numbers appear multiple times in the question or the code, where a strategy for grounding is needed. Let $C_{q}(q_{i})$ be the count of $q_{i}$ in the question, and let $C_{c}(c_{i})$ be the count of $c_{i}$ in the code. The strategy will be broken down into several scenarios:
  1. 1.
    If there is a number $q_{i}$, such that $C_{q}(q_{i})<C_{c}(q_{i})$, then the question will be discarded for further problem generations. The reason to discard this kind of problem is that the relation between the numbers can be very convoluted. One number in the question might correspond to zero, one, or many numbers in the code, and finding the correct correspondence requires human evaluation.
  2. 2.
    There is a number $q_{i}$, such that $C_{q}(q_{i})>1$ and $C_{q}(q_{i})=C_{c}(q_{i})$. Similar to the previous scenario, one number in the question can correspond to zero, one, or many numbers in the code. However, after some manual inspection, we find that most questions under this scenario have a one-to-one correspondence between the numbers in the questions and the codes. Therefore, we have devised a method to ensure that the right correspondence of numbers can be retrieved. Given that the number appears $C_{q}(q_{i})$ times in the question and the code, we constructed $C_{q}(q_{i})!$ pairs of one-to-one correspondences to the numbers from the syntax tree to the question. Then, random numbers are generated to combine the equivalences of correspondences. For example, given that the question is "Mary has 5 apples and 5 oranges, how many fruits does Mary have?". Although the two "5"s correspond to different representations in the question, due to the communicative nature of addition, the two "5"s are interchangeable in the syntax tree. Thus, the two pairs of correspondence for this question can be combined, and there is no ambiguity in the grounding. After the correspondences are combined, one valid sequence of numbers for the question is randomly generated for the code that ensures that all correspondences of the syntax tree output different answers. Then, this version of the question is asked to GPT-4 to obtain the answer to the question. The correspondence that has the same answer will be used, and if no correspondence has the correct answer, the question will be discarded. An example of this type of question is: "The pizzeria sells small pizzas for \$2 and large pizzas for \$8. They sold \$40 in pizzas. If they sold 8 small pizzas, how many large pizzas did they sell?". One valid and distinguishing sequence of numbers for the question is $(2,3,31,5)$, which will result in different answers $5$ and $7$.
  3. 3.
    There is a number $c_{i}$, such that $C_{q}(c_{i})>C_{c}(c_{i})$. Similar to the reasoning for the first case, these kinds of questions are eliminated for problem generations.
Appendix G Feature Analysis
Features MetaMath-7B-V1.0 vicuna-13b-v1.5 CodeLlama-34b-hf gpt-3.5-turbo-1106
Addition Count -0.0032 0.0080 -0.0146 0.0098
Divide Count -0.0648 -0.1137 0.0142 -0.0804
Minus Count -0.0187 0.0040 -0.0254 -0.0195
Multiply Count -0.0328 0.0214 0.0031 -0.0520
Constant Count 0.0923 0.0782 0.0027 0.1559
Variable [8, 32) 0.0011 -0.0142 -0.0100 0.0117
Answer [2, 8) 0.2215 0.1377 0.0632 -0.0736
Answer [8, 32) 0.2437 0.1104 0.0215 -0.0787
Answer [32, 128) 0.2610 0.1670 0.0183 -0.0508
Answer [128, 512) 0.2267 0.0998 -0.0259 -0.0726
Answer [512, 2048) 0.0076 0.0864 -0.0380 -0.0343
Answer [2048, 8192) -0.1987 -0.0775 -0.0434 0.0336
Convert to Int 0.2400 0.1664 0.0470 0.2476
Operation Count -0.1196 -0.0804 -0.0227 -0.1421
Variable Count 0.0722 0.0254 0.0074 0.0939
Constant 0.2840 0.1840 0.0328 0.3919
Features Llama-2-13b-hf MetaMath-70B-V1.0 Mistral-7B-v0.1 WizardMath-13B-V1.0
Addition Count -0.0163 -0.0287 0.0102 -0.0118
Divide Count 0.0523 0.0135 0.0134 -0.0423
Minus Count -0.0152 -0.0271 0.0097 -0.0137
Multiply Count -0.0248 -0.0473 -0.0422 -0.0369
Constant Count -0.0051 0.1136 0.0478 0.0958
Variable [8, 32) -0.0349 0.0286 -0.0383 0.0266
Answer [2, 8) 0.0205 0.0588 0.1006 0.1448
Answer [8, 32) 0.0256 0.0930 0.1065 0.1199
Answer [32, 128) -0.0011 0.1186 0.0479 0.1234
Answer [128, 512) -0.0006 0.1506 0.0325 0.0964
Answer [512, 2048) -0.0186 0.0771 -0.0106 -0.0637
Answer [2048, 8192) -0.0242 0.0105 -0.0008 -0.1885
Convert to Int -0.0435 0.1771 -0.0894 0.0498
Operation Count -0.0040 -0.0895 -0.0089 -0.1047
Variable Count 0.0217 0.1068 0.0238 0.0795
Constant 0.0205 0.3100 0.0805 0.2800
Table 7: Coefficients for some selected features after fitting linear regression models to the features of problems and the correctness of the eight models. Note that coefficients should not be compared across models, only with other features within the same model. For example, we can observe that MetaMath-7B performs well when the answer is in the range [2, 512), while performance declines when the answer is in the range [512, 8192). The observation matches the accuracy shown in Figure 7.
Figure 7: The accuracy of eight models is plotted against four relevant features when tasked with all questions generated by method M3. The top left graph shows the accuracy of models against problems with various counts of multiplications, and the top right graph compares problems with various counts of all operations, including binary and unary. The middle graph shows the accuracy of models with different ranges of generated numbers, and the bottom graph shows the accuracy with different ranges of final answers.

### v2

@@
 We identify shared vulnerabilities among LLMs and propose a cost-effective approach to attack high-cost models.
-Additionally, we conduct automatic analysis on math problems and investigate the cause of failure to guide future research on LLM’s mathematical capability.
-[^1] Data and code are available: https://github.com/ruoyuxie/adversarial_mwps_generation
-LLM-Resistant Math Word Problem Generation via Adversarial Attacks
-Roy Xie      Chengxuan Huang      Junlin Wang      Bhuwan Dhingra Duke University {ruoyu.xie, jonathan.huang, junlin.wang2}@duke.edu {bdhingra}@cs.duke.edu
+Additionally, we conduct automatic analysis on math problems and investigate the cause of failure, offering a nuanced view into model’s limitation.
 1 Introduction
@@
 At the same time, adversarial attacks on LLMs have gained more attention due to increased awareness of the potential risks associated with LLMs.
-Most work on adversarial attacks focuses on developing prompts to elicit specific outputs from LLMs (Wang et al., 2021; Zhu et al., 2023).
-Recent work suggests that even the most powerful aligned LLMs are vulnerable to such attacks (Zou et al., 2023; Carlini et al., 2023).
-Hence, we might expect that as LLMs become stronger, detecting their outputs will become more difficult, but adversarial examples may still persist Ilyas et al.
-(2019).
+Most work on adversarial attacks focuses on developing prompts to elicit specific outputs from LLMs (Zhang et al., 2020; Zou et al., 2023; Carlini et al., 2023) Recent work suggests that even the most powerful aligned LLMs are vulnerable to such attacks (Zou et al., 2023; Carlini et al., 2023).
+Hence, we might expect that as LLMs become stronger, detecting their outputs will become more difficult, but adversarial examples may still persist (Ilyas et al., 2019; Wei et al., 2023).
 To this end, we introduce a new paradigm for generating homework assignments that LLMs cannot solve by utilizing adversarial attacks.
 We focus on math word problems (MWPs), which present a unique and challenging intersection of language and mathematics.
-In this work, we ask the question: “Can LLMs still solve an MWP after changing its numeric values?” We aim to create MWPs that LLMs are unable to solve while maintaining the coherence and original difficulty of the problems.
-This goal diverges from typical adversarial attacks in machine learning, where small, often imperceptible changes are made to input data (images, text, etc.) to deceive models into making incorrect predictions.
-Here we do not focus on making imperceptible changes, but rather on preserving the coherence, style, and, most importantly, difficulty of the original problems.
+In this work, we ask the question: “Can LLMs still solve an MWP after changing its numeric values?” We aim to create MWPs that LLMs are unable to solve while maintaining the original difficulty and coherence of the problems.
 Our aim is not just to challenge LLMs but to do so in a way that reflects real-world educational standards to ensure the MWPs remain educationally valuable and relevant.
-We generate adversarial examples by editing the numeric values in MWPs.
-While conceptually simple, doing this effectively is challenging for a couple of reasons: (i) the answer of the resulting problem changes and must be recomputed to assess the robustness of LLMs; and (ii) preserving the difficulty and validity of the modified problem requires us to ensure that all the intermediate calculations in the new problem are consistent with the original problem.
-For example, if the original problem did not have fractions as intermediate values, in an educational context we would not want the modified problem to have such values either.
-To tackle these challenges, we first convert MWPs to a code representation and leverage abstract syntax trees (ASTs) to systematically generate adversarial problems.
-We evaluate several major LLMs on their math-solving ability under adversarial settings, demonstrating the effectiveness of our method with a significant degradation across various LLMs.
-We investigate universal attacks and attack transferability, proposing a cost-effective approach to attack high-cost models, achieving a 90% request reduction rate without sacrificing performance.
-We verify the validity of our generated problems by conducting human evaluations.
-Lastly, we conduct in-depth automatic analyses to investigate the mathematical vulnerabilities in LLMs and find, unsurprisingly, that the values of the answer and the number of operations have strong correlations with correctness.
-Figure 1: Method Overview: given a MWP that an LLM can correctly solve, our method first transforms it into Python code.
-Next, the Python code is converted into an AST representation which is used to generate adversarial problems by modifying the numeric values in the problem.
-We place constraints on the nodes of the AST representation to ensure that the modified problem maintains the same difficulty level as the original problem.
+Figure 1: Method Overview Given a MWP that an LLM can correctly solve, our method first transforms it into Python code.
+The Python code then is converted into an AST representation which is used to generate adversarial problems by modifying the numeric values in the problem.
+We place constraints on the nodes of the AST to ensure that the modified problem maintains the same difficulty level as the original problem.
 Despite this, we find that the resulting adversarial examples cause LLMs to predict incorrect answers.
+While conceptually simple, doing this automatically is challenging as directly modifying numbers in the problem could lead to nonsensical problems.
+For example, consider the problem “A class has 6 male students and half as many female students.
+How many students in total?” Directly change the number of male students from 6 to 5 (or any odd numbers) without checking its intermediate computational steps might result in “2.5 female students,” which is not only illogical but also introducing fraction to the problem, which might change the indented difficulty.
+Similarly, changing the number from 6 to 624 might make the problem unrealistic and much harder for students.
+To generate meaningful and coherent problem variations, it is essential to consider the logical implications of number modifications and ensure that the resulting problem remains plausible and solvable.
+We further discuss the importance of educational constraints in §3.2.
+To effectively generate modifications in large scale to assess the robustness of LLMs also requires: (i) The answer to the altered problem changes must be recomputed automatically; and (ii) Preserving the difficulty and validity of the modified problem requires us to ensure that all the intermediate calculations in the new problem are consistent with the original problem.
+To tackle these challenges, we first convert MWPs to a code representation and leverage abstract syntax trees (ASTs) to map each calculation step into a node.
+We then define educational constraints for each node to ensure all the desired properties are preserved for the generated new problem.
+In this work, we evaluate several major LLMs and demonstrate the effectiveness of our method by achieving a significant attack success rate (ASR).
+Our approach outperforms the previous rephrasing attack by an average of 60% ASR.
+We investigate universal attacks and attack transferability, proposing a cost-effective approach to attack high-cost models (e.g., GPT-4) by reducing API request calls by 90% without sacrificing performance.
+We conduct human evaluations to verify that our generated problems indeed preserve the original coherence and difficulty.
+Furthermore, we perform a regression analysis and find that our adversarial examples exploit different weaknesses of each model, offering valuable insights and a nuanced view into model’s limitation.
 2 Background and Related Work
 Fair Evaluation for Educational Purpose
-The increasing sophistication of LLMs has made it challenging for plagiarism detection tools to accurately identify content generated by machines (Chaka, 2023; Liang et al., 2023).
-Previous work suggests that detecting LLM-generated content may become nearly impossible as LLMs improve (Sadasivan et al., 2023).
-The issue extends beyond text-based content; various other disciplines are also threatened by the advancing capabilities of LLMs.
-For instance, LLMs can now provide step-by-step solutions for MWPs, which complicates the matter of fair evaluation, as students may pass off machine-generated solutions as their own without understanding the underlying mathematical concepts.
-It remains an open question of how to ensure fair evaluation for students in the presence of LLMs.
+As LLMs become more adept at generating human-like text, it becomes increasingly difficult to distinguish between student-and machine-generated content (Chaka, 2023; Liang et al., 2023), which poses significant challenges for educational institutions in ensuring fair evaluation of student work (Yan et al., 2024).
+This issue also extends beyond traditional written assignments, as LLMs can now provide detailed solutions to complex problems across various disciplines (Abedi et al., 2023).
+Consequently, educators must develop new strategies to assess student understanding and maintain the integrity of the evaluation process (Liu et al., 2023).
 LLMs Math-solving Ability
-The capability of LLMs to solve math problems is currently a hot topic (OpenAI, 2023; Yu et al., 2023; Xu et al., 2023).
-Studies found that LLMs can significantly improve their math-solving ability through prompt engineering (Yu et al., 2023; Imani et al., 2023), such as chain-of-thought (CoT) (Wei et al., 2022).
+Our work also relates closely to LLMs math reasoning ability (OpenAI, 2023; Yu et al., 2023; Xu et al., 2023).
+Studies found that LLMs can significantly improve their math-solving ability through prompt engineering (Yu et al., 2023; Imani et al., 2023a), such as chain-of-thought (CoT) (Wei et al., 2022).
 Converting the MWPs’ solving steps into code or symbolic representations can also improves LLM’s performance (Li et al., 2023; He-Yueya et al., 2023; Gao et al., 2023).
-While different prompting strategies exist, our focus is to generate adversarial examples that LLMs are unable to solve.
-Therefore, we aim to minimize prompt engineering by using a unified prompt with zero-shot CoT for all queries in this work.
-[^2] Prompt template is in Appendix A.2.
+[^1] While different prompting strategies exist, our focus is to generate adversarial examples that LLMs are unable to solve.
+Therefore, we try to minimize prompt engineering by using a unified prompt with zero-shot CoT for all evaluations.
+Prompt can be found in Appendix A.2.
 Adversarial Attacks on MWPs
-Adversarial attacks on LLMs focus on modifying prompts to cause LLMs to output desired content (Zhang et al., 2020; Zou et al., 2023; Carlini et al., 2023).
-Similarly, our focus is to modify the MWPs to cause LLMs to output incorrect answers.
+Adversarial attacks on LLMs involve modifying prompts to alter their behavior (Zhang et al., 2020; Zou et al., 2023; Carlini et al., 2023).
+Similarly, we modify MWPs to cause LLMs to output incorrect answers.
 Bubeck et al.
-(2023) conducts a limited number of memorization tests on MWPs by randomly changing the numeric values between -10 and 10, suggesting that state-of-the-art LLMs do not simply rely on memorizing problems but on applying a general solution method.
-However, the number of examples in Bubeck et al.
-(2023) is relatively small to reach such a conclusion, and we will show in our work that modifying numbers in MWPs does cause LLMs to fail on a larger scale.
-Zhou et al.
-(2023), on the other hand, rephrases MWPs by changing the words within the problem and keeping all the numeric values the same.
-However, this approach makes it challenging to preserve the problem’s original context and structure, and sometimes could introduce inconsistent and problematic content, as shown in Appendix B.
-Additionally, Zhou et al.
-(2023) also requires extensive human validation to ensure modification quality.
-Therefore, we focus on changing numerical values, which ensures the preservation of the underlying logic.
+(2023) conducted limited memorization tests on MWPs by randomly changing numeric values between -10 and 10, suggesting that state-of-the-art LLMs do not solely rely on memorization but apply general solution methods.
+However, their study’s sample size was relatively small, and we demonstrate that modifying numbers in MWPs causes LLMs to fail on a larger scale.
+On the other hand, Zhou et al.
+(2023) rephrased MWPs by changing words while preserving numeric values.
+This approach risks altering the original context and introducing inconsistencies and problematic content (see Appendix B), necessitating extensive human validation.
+Therefore, we focus on changing numerical values, preserving the underlying logic.
 To the best of our knowledge, this is the first study to undertake such an endeavor.
@@
 An overview of our method can be found in Figure 1.
-3.1 Preliminaries
 We denote an original problem-answer pair as $p=(x,y)$, where $x$ is a sequence of tokens $x_{1},\dots,x_{n}$ in the problem and $y$ is its ground-truth answer.
+We define $G$ as a ground-truth function that computes the correct answer for a math problem [^2] In our case $G$ is generated Python code.
+and $F$ as a function which maps the elements of a sequence as: $F(\tilde{x}_{i})\sim\mathbb{R}$ (i.e., a random real number) if $x_{i}$ is numeric and $F(\tilde{x_{i}})=x_{i}$ otherwise.
 We denote the set of all possible adversarial modifications to it as:
 $A(x,y)=\{(\tilde{x},\tilde{y}):\tilde{x}_{i}\in F(x),\tilde{y}=G(\tilde{x})\}.$ (1)
-Here $G$ is a ground-truth function that computes the correct answer for a math problem (in our case this is generated Python code).
-$F$ is a function which maps the elements of a sequence as: $F(\tilde{x}_{i})\sim\mathbb{R}$ (i.e., a random real number) if $x_{i}$ is numeric and $F(\tilde{x_{i}})=x_{i}$ otherwise.
-This entire set of adversarial examples will also consist of many unnatural and difficult problems, hence later in §3.3 we will introduce filtering constraints for selecting the adversarial examples that preserve difficulty.
+$A(x,y)$ will also consist of many unnatural and difficult problems, hence later in §3.2 we will introduce filtering constraints for selecting the adversarial examples that preserve difficulty.
 3.1 Mapping From MWPs to Code to Tree
-To structurally modify MWPs, we first convert MWPs into Python code by prompting GPT-4-turbo to generate the code based on the problem-solving steps provided by the datasets.
-[^3] Prompt can be found in Appendix A.1.
-As a sanity check, we first select 200 MWPs each from GSM8K (Cobbe et al., 2021) and MultiArith (Roy and Roth, 2015) and check if the generated program results in correct answers.
-Given the execution steps in code are almost always deterministic, encouragingly, we found only 2 and 5 generated codes have incorrect logic or answers in GSM8K and MultiArith, respectively.
-We further investigate the generated Python code in §5.3.
-Next, we utilize AST, a tree representation for the abstract syntactic structure of programming language code, to convert the generated Python code into a tree.
+To structurally modify MWPs, we first use GPT-4 to generate the Python code that reflects the solution steps given the problem and its final answer.
+[^3] We only require the problem and its final answer, which is almost always given in the context of math problems.
+Prompt can be found in Appendix A.1.
+As a sanity check, we select 200 MWPs each from GSM8K (Cobbe et al., 2021) and MultiArith (Roy & Roth, 2015) and check if the generated program results in correct answers.
+Given the execution steps in code are almost always deterministic, encouragingly, we found only 1% and 2.5% generated codes have incorrect logic or answers in GSM8K and MultiArith, respectively.
+We further investigate the generated Python code in §5.
+Next, we utilize AST, a tree for the abstract syntactic structure of programming language code, to convert the generated Python code into a tree representation.
 We build the ASTs by traversing through the Python code and constructing nodes corresponding to each statement.
 The final print statement that outputs the answer is the root of the tree.
-There are mainly two types of nodes: operation nodes, which carry out the operations in the tree and are the non-leaf nodes, and variable nodes, which correspond to the numeric values from the problem and are the leaf nodes.
-We discuss the nodes in detail in Appendix C.
+Each AST has mainly two types of nodes: operation nodes, which carry out the operations in the tree and are the non-leaf nodes, and variable nodes, which correspond to the numeric values from the problem and are the leaf nodes.
+We discuss the nodes in AST in detail in Appendix C.
 3.2 Adversarial Example Generation
-In an adversarial setting, attackers can perform multiple attacks on the system. Similarly, we generate multiple adversarial examples which are different versions of the original MWPs to attack LLMs.
+In an adversarial setting, attackers can perform multiple attacks on the system, and any successful attack signifies the possible flaws of the system (Carlini & Wagner, 2017).
+Similarly, we generate adversarial examples which are different versions of the original MWPs to attack LLMs.
 Educational Context
@@
 For instance, changing one-digit multiplication to four-digit multiplication significantly alters the problem’s complexity.
-Furthermore, to keep the mathematical logic intact, the order of magnitude of numbers in the problem should remain unchanged.
-For example, altering a problem from “subtracting 2 apples from 5” to “subtracting 4 apples from 2” introduces the concept of negative numbers, potentially misaligning with the original problem’s context.
-Lastly, changes should be contextually coherent.
-For instance, a generated problem with “the first half of 3 people” would be confusing.
+To maintain mathematical logic intact, the order of magnitude of numbers in the problem should also remain unchanged.
+For example, changing a problem from “Jack ate 2 out of a total of 5 apples” to “Jack ate 5 out of a total of 2 apples” not only introduces the concept of negative numbers but also alters the problem’s logical structure and coherency.
+Furthermore, such modifications can lead to unrealistic scenarios that deviate from the intended learning objectives and may cause confusion.
 Filtering Constraints
 To minimize the difference between the original MWP and its adversarial counterpart while adhering to the educational context, we define a set of Boolean constraints for each node to control the generation quality.
-Given the original node value $h$ and the new value of node $h^{\prime}$, a newly generated problem is valid if and only if the $h^{\prime}$ satisfies the same set of constraints that $h$ has, for all nodes.
+Given the original node value $h$ and its new value $h^{\prime}$, a newly generated problem is valid if and only if $h^{\prime}$ satisfies the same set of constraints that $h$ has, for all nodes.
 A list of node constraints is:
@@
 Constrictive Generation Methods
-The node constraints only ensure that the adversarial examples are valid in the numerical sense; instances that don’t make logical sense like “$28$-hour work day”, which is not common, can still be generated.
-Furthermore, different values for the variable nodes can lead to vastly different difficulty levels, as $h^{\prime}$ can be significantly larger than $h$.
-To mitigate this, we propose three generation methods to distinguish generated adversarial examples with varying levels of difficulty.
+The node constraints ensure the adversarial examples are valid in the numerical sense; instances that don’t make logical sense like “$28$-hour work day”, which is not common, can still be generated.
+Different values for the variable nodes can also lead to vastly different difficulty levels, as $h^{\prime}$ can be significantly larger than $h$.
+To mitigate this, we propose three generation methods to distinguish generated adversarial examples with different levels of difficulty.
 We decompose the values $h\approx a\times 10^{b}$, where $a$ is an integer that is not divisible by 10 and with at most $c$ digits.
-Depending to the numbers $h$, $a$, and $b$, the generation methods are defined as follows:
-  • $M1$: Free Generation: regardless of the value of $a$, $a^{\prime}$ is drawn from a uniform distribution between $1$ and $10^{c}$.
+Depending on the numbers $h$, $a$, and $b$, the generation methods are defined as follows:
+  • $M1$: Free Generation: Allowing the generated problem to have a wide range of numbers with minimal constraints, with each $h^{\prime}$ to be $a^{\prime}\times 10^{b}$.
+Regardless of the value of $a$, $a^{\prime}$ is drawn from a uniform distribution between $1$ and $10^{c}$.
 If the variable node $v$ is a divisor of a division node, then $a^{\prime}$ is drawn from a uniform distribution between $1$ and $\left\lfloor 10^{c/2}\right\rfloor$.
-The newly generated value $h^{\prime}$ will be $a^{\prime}\times 10^{b}$.
-The reason for picking smaller numbers for a divisor is that the probability of picking a number from $1$ and $10^{c}$ that is divisible for another number from $1$ and $10^{c}$ is very low for big values of $c$.
-  • $M2$: Count of Digits: this generation method ensures that the newly generated value $h^{\prime}$ has the same number of digits as $h$.
+  • $M2$: Count of Digits: Constraining the generated value $h^{\prime}$ has the same number of digits as $h$.
 For example, for $h=100$, some possible $h^{\prime}$ are 942, 589, or 264.
@@
 For more variability, $h^{\prime}$ can range from 1 to 99 if $h$ has only one digit.
-For decimal variable values like $h=1.25$, we will use $a=125$ to generate numbers like $a^{\prime}=473$, then convert them back to $h^{\prime}=4.73$.
-  • $M3$: Count of Scientific Numbers: this generation method tries to ensure that the new value shares a similar scientific digit count and is within a similar range to the original value.
-For example, for $h=1500$, $h^{\prime}$ can be 1700, 800, 1200.
-The rationale is that changing a number to a larger one does not necessarily make calculations harder, but more scientific numbers do.
+For decimal variable values like $h=1.25$, we use $a=125$ to generate numbers like $a^{\prime}=473$, then convert them back to $h^{\prime}=4.73$.
+  • $M3$: Count of Scientific Numbers: Constraining the new value shares a similar scientific digit count and is within a similar range to the original value.
+For example, for $h=1500$, $h^{\prime}\sim\text{Pois}(h)$ can be 1700, 800, 1200, where Pois is the Poisson distribution.
+The rationale is that larger numbers do not result in more difficult problems, but more scientific numbers do.
 An example would be comparing the numbers 150,000 and 172,568.
-$M3$ is the most restrictive generation, followed by $M2$, and $M1$ is the least restrictive, therefore:
+$M3$ is the most restrictive generation, followed by $M2$, and $M1$ is the least restrictive:
 $M1(A(x,y))\supseteq M2(A(x,y))\supseteq M3(A(x,y)).$ (2)
-The more restrictive a method is, the closer the difficulty levels and coherence are between the adversarial examples and the original problem.
-Therefore, we use $M3$ as our main generation method as it ensures the adversarial examples comply with the constraints while also gauging the difficulty of problems.
+The more restrictive a method is, the closer the difficulty levels and coherence are between the an adversarial example and its original version.
+Therefore, we use $M3$ as our main generation method as it ensures the adversarial examples comply with all constraints to maintain the original difficulty and coherence.
 Although generations from $M1$ and $M2$ may not fit in the educational context, we are interested in studying LLMs’ math-solving abilities by simply changing the numeric values.
 We discuss how different methods impact model performance in §4.2.
-We describe each generation method in more detail in Appendix D and present some sample generations based on each method in Table 5.
+We describe each generation method in more detail in Appendix D and present the generated examples based on three methods in Table 5.
 4 Experiments and Results
@@
 Datasets
-We generate adversarial examples from GSM8K and MultiArith.
+We generate problem variants from GSM8K (Cobbe et al., 2021) and MultiArith (Roy & Roth, 2015).
 Both datasets are commonly used for evaluating LLMs’ mathematical capabilities, and MultiArith is sourced from math worksheets used by elementary school students to practice math problems (Roy et al., 2015).
+[^4] While MWPs exhibit different levels of difficulty, we focus on MWPs at this particular difficulty level as a proof of concept in this work.
 Models
-We conduct experiments on 7 open-source models: MetaMath 7B, 70B (Yu et al., 2023), Mistral 7B (Jiang et al., 2023), Llama-2 13B (Touvron et al., 2023), WizardMath 13B (Xu et al., 2023), Vicuna 13B (Chiang et al., 2023), CodeLlama (Roziere et al., 2023), and 2 closed-source models: GPT-4-Turbo (OpenAI, 2023) and GPT-3.5-Turbo (OpenAI, 2022).
+We conduct experiments on 7 open-source models: MetaMath 7B, 70B (Yu et al., 2023), Mistral 7B (Jiang et al., 2023), Llama 2 13B (Touvron et al., 2023), WizardMath 13B (Xu et al., 2023), Vicuna 13B (Chiang et al., 2023), CodeLlama (Roziere et al., 2023), and 2 closed-source models: GPT-4 (OpenAI, 2023) and GPT-3.5 (OpenAI, 2022).
 [^5] Specifically, we use GPT-4-0125-preview and GPT-3.5-turbo-0125.
@@
 The model selection was largely based on the LLMs’ math-solving ability.
-Our intention is to evaluate a wide range of LLM performances, including the math-tuned LLMs such as MetaMath and WizardMath, popular open-source LLMs such as Llama-2 and Mistral, and the state-of-the-art API-based GPT models.
+Our intention is to evaluate a wide range of LLM performances, including the math-tuned LLMs such as MetaMath and WizardMath, popular open-source LLMs such as Llama 2 and Mistral, and the API-based GPT models.
 Metrics
-Given a set of original problem-answer pairs $P$ and a LLM $L$, we use the follwing evaluation metrics:
-  • Original Accuracy (OA): the accuracy of $L$ on the original problems:
+We follow the previous adversarial attack literature to measure the attacks, which looks for at least one perturbation that fools the model (Croce et al., 2020; Pruthi et al., 2019; Jia & Liang, 2017).
+A problem is considered incorrect if it has at least one incorrect variation.
+Such variation is also known as an adversarial example.
+Given a set of original problem-answer pairs $P$ and a LLM $L$, we define: [^6] OA, AA, ASR are reported in % in this work.
+  • Original Accuracy (OA): the accuracy of $L$ on the original problems,
     $\text{OA}(L)=\frac{\sum_{(x,y)\in P}\mathbf{1}\{L(x)=y\}}{|P|}.$ (3)
-  • Attack Accuracy (AA): given an indicator function $I_{xy}$:
-    $I_{xy}=\mathrm{1}\left[\forall(\tilde{x},\tilde{y})\in A(x,y):L(\tilde{x})=\tilde{y}\right],$ (4)
-    $L$’s accuracy on the adversarial examples:
-    $\text{AA}(L)=\frac{\sum_{(x,y)\in P}I_{xy}}{|P|}.$ (5)
-  • Attack Success Rate (ASR): the relative decrease in accuracy due to adversarial modifications:
+  • Attack Accuracy (AA): the accuracy of $L$ on the adversarial modifications,
+    $\text{AA}(L)=\frac{\sum_{(x,y)\in P}\mathrm{1}\left[\forall(\tilde{x},\tilde{y})\in A(x,y):L(\tilde{x})=\tilde{y}\right]}{|P|}.$ (4)
+  • Attack Success Rate (ASR): relative decrease in accuracy due to adversarial modifications,
     $\text{ASR}(L)=\frac{\text{OA}(L)-\text{AA}(L)}{\text{OA}(L)}.$ (5)
-4.2 Rephrasing Attacks
-We examine the effectiveness of a commonly used rephrasing approach, where the modifications are centered around the words within the MWPs.
-Zhou et al.
-(2023) freeze the logical entities (e.g., entity names and values) and iteratively go through each unfrozen token to swap it with similar tokens that could cause LLMs to fail.
-They then validate the adversarial example with human manual checking to ensure the example’s quality.
-Zhou et al.
-(2023) present the RobustMath dataset, which includes 214 original problems from a combination of GSM8K and MultiArith, and generate 300 rephrased versions.
-As a baseline, we report their ASR in Table 2.
-We observe that rephrasing attacks result in lower accuracy for 4 out of the 7 models; however, some models show improved accuracy, suggesting that such attacks may not generalize across different LLMs.
 MultiArith GSM8K
@@
 MetaMath 7B 100 74 26 10 90 0 100 95 28 71 9 91 0 100
+WizardMath 13B 89 20 78 5 94 0 100 89 11 88 2 98 0 100
+Vicuna 13B 76 4 95 1 99 0 100 60 0 100 0 100 0 100
 Llama 2 13B 12 0 100 0 100 0 100 10 0 100 0 100 0 100
-WizardMath 13B 89 20 77 5 94 0 100 89 11 88 2 98 0 100
-Vicuna 13B 76 4 95 1 99 0 100 60 0 100 0 100 0 100
 CodeLlama 34B 11 0 100 0 100 0 100 6 0 100 0 100 0 100
 MetaMath 70B 99 86 13 30 70 0 100 98 50 49 17 83 0 100
-GPT-3.5-turbo 97 74 24 47 52 0 100 91 52 43 31 66 0 100
-Table 1: Main Results: we generate adversarial examples based on the given constraint.
-Overall, simply changing the numeric values in the original math problems causes a consistent performance drop across all models, even with the most restrictive generation method.
-4.3 Number of Adversarial Examples
-We investigate how the number of adversarial examples impacts model performance. Figure 2 shows the average number of possible adversarial examples per question for each constraint.
-Figure 2: Possible Generations: with 30,000 attempts for each given constraint, we calculate the average number of adversarial examples generated for each problem.
-Figure 3: Increase Examples: we report the model performance given various number of adversarial examples. For several models, only 10 adversarial examples are enough to degrade the performance.
-While on average each problem generates over a thousand modifications, it is worth noting that the actual number of possible generations varies significantly based on the values and the number of variable nodes in a problem.
-For example, a problem with only one digit value and one variable node has a potential maximum of 9 possible modifications under the $M3$ generation method.
-Given this reason and computational cost considerations, in this work we randomly select 100 problems from MultiArith and GSM8K separately and generated 100 adversarial examples for each problem.
-Despite the seemingly small number of problems, the creation of 30,000 adversarial examples per dataset given three different generation methods (totaling 60,000 examples across both datasets) is significant.
-We present the GSM8K results in Figure 3, observing a clear pattern of accuracy decline across all models as the number of examples increases.
-Interestingly, for several models, just 10 adversarial examples are enough to degrade the performance.
-4.4 Main Results
-We evaluate each adversarial example for each problem against all models given different generation methods.
-A problem is considered incorrect if there exists at least one adversarial example under that problem that results in an incorrect answer.
-We refer such problem as successful attack.
-Here we do not run the full generated adversarial examples against GPT-4 due to cost consideration.
-Instead, we propose a cost-effective way to query expensive models in §4.6.
-We report the model performance with all three generation methods in Table 1.
-Ranging from the most lenient method, $M1$, where the numbers could be fairly wild, to the most stringent condition, $M3$, we observe a consistent accuracy drop across all models, even for the strong models (e.g., math-tuned, larger size, or API-based).
-For the most original-problem-like generation, $M3$, we observe that the weaker models (e.g., smaller size or general-purpose) fail to generate any correct answers in every case.
-As a comparison to rephrasing attack, we present the average ASR of $M3$ in Table 2.
-Even as the most restrictive generation method, $M3$ significantly outperforms RobustMath in every model.
-Model ASR (RobustMath) ASR (ours)
-Mistral 7B 0.0 100.0
-MetaMath 7B 13.0 48.5
-Llama-2 13B 0.0 100.0
-WizardMath 13B 1.0 82.5
-Vicuna 13B 0.0 97.5
-CodeLlama 34B 67.1 100.0
-MetaMath 70B 11.1 31.0
-GPT-3.5-Turbo 16.9 33.5
-Table 2: Comparison Result: given that RobustMath comprises MultiArith and GSM8K, we calculat the average ASR of $M3$ from our results for MultiArith and GSM8K.
-Our method significantly outperforms RobustMath in every model.
-4.5 Universal Attacks
-We are interested in whether there exist problems that all LLMs fail, also known as universal attacks.
-We count the number of successful attacks from each model and calculated the percentage of common successful attacks among them.
-We present the results in Table 3 and observe a clear pattern of decreasing universal attacks with an increased number of models.
-This suggests that diversity among models can lead to increased resilience against such attacks.
-We also observe that the impact of universal attacks varies significantly across different generation methods, with $M3$ having the lowest number of universal attacks.
+GPT-3.5 97 74 24 47 52 0 100 91 52 43 31 66 0 100
+Average 65.1 32.3 66.9 11.6 88.1 0 100 59.8 17.6 81.4 7.4 92.3 0 100
+Table 1: Main Attacks We generate MWPs based on different generation methods.
+Overall, simply changing the numeric values in the problems causes a consistent performance drop across all LLMs, even with the most restrictive generation method.
+Baseline
+As a baseline, we examine the effectiveness of a commonly used rephrasing approach, where the modifications are centered around the words within the MWPs.
+Zhou et al.
+(2023) freeze the logical entities and iteratively swap each unfrozen token with similar tokens that cause LLMs to fail.
+They then validate the adversarial example with human manual checking.
+Zhou et al.
+(2023) created the RobustMath dataset, which includes 214 original problems from a combination of GSM8K and MultiArith with their 300 rephrased versions.
+We report the results in Appendix E, observing ASR for 4 out of the 7 models; however, some models even show improved accuracy given these adversarial examples, suggesting that such attacks may not generalize well across different LLMs.
+4.2 Our Attacks
+Problem Variation Generation
+We set the number of attempts to generate problem variants to be 30,000 and present the average number of generated variants per problem for each $M$ in Appendix G.
+While on average each problem generates over a thousand variants, it is worth noting that the actual number of possible generations varies based on the original values and the number of variable nodes in a problem.
+[^7] For example, a problem with only one digit value and one variable node has a potential maximum of 9 possible modifications under $M3$.
+Given this reason and cost considerations, in this work we equally select 100 random problems from MultiArith and GSM8K and generate 100 variants for each selected problem in each $M$.
+Despite the seemingly small number of selected problems, the total data points are up to 60,000.8883 generation methods $\times$ 200 selected problems $\times$ 100 variants = 60,000.
+Main Attacks
+We report the model performance from all $M$s in Table 1.999We do not run the full dataset against GPT-4 due to cost considerations.
+Instead, we propose a cost-effective approach to query expensive models in Efficient Attack section.
+Ranging from the most lenient method, $M1$, where the numbers could be fairly wild, to the most stringent one, $M3$, we observe a consistent accuracy drop across all models, even for the strong models (e.g., math-tuned, larger size, or API-based).
+For the most original-problem-like generation, $M3$, we observe that weak models (e.g., smaller size or general-purpose) fail to generate any correct answers in every case.
+The left of Table 2 shows each model’s average percentage of incorrect variants.
+It is worth noting that these percentages are highly correlated with the reported ASR, with GPT-3.5 and MetaMath being the most robust, and CodeLLama and LLama 2 being the least robust models.
+Furthermore, we compare $M3$ with rephrasing attacks in Table 3.
+Our method significantly outperforms the baseline in every model, resulting in 60.5 ASR point improvement on average.
+We also analyze how ASR changes given different numbers of attacks in Appendix H.
+Interestingly, for several models, just 10 attacks are enough to degrade the performance significantly.
+Models
+AE. (%)
+Mistral 7B 70.5
+MetaMath 7B 15.2
+Llama 2 13B 91.6
+WizardMath 13B 21.0
+Vicuna 13B 50.3
+CodeLlama 34B 91.8
+MetaMath 70B 7.9
+GPT-3.5 6.9
 Model Ct. M3 (%) M2 (%) M1 (%)
@@
 MetaMath 7B 2 70.0 90.0 100.0
-Llama-2 13B 3 67.0 90.0 100.0
+Llama 2 13B 3 67.0 90.0 100.0
 WizardMath 13B 4 52.0 80.0 100.0
@@
 MetaMath 70B 7 9.0 49.0 87.0
-GPT-3.5-Turbo 8 9.0 46.0 83.0
-Table 3: Universal Attack: the number of universal attacks decreases as the number of models increases, implying that diversity among models enhances resilience against such attacks.
-4.6 Efficient Attacks
-Figure 3 shows that scaling up the number of adversarial examples results in a steady performance drop for all models.
-However, this scaling might not be viable for API-based models such as GPT-4 or GPT-3.5 due to request limits and high costs.
+GPT-3.5 8 9.0 46.0 83.0
+Table 2: Incorrect Variants (Left) The average percentage of incorrect variants per problem for each model.
+A detailed count distributions can be found in Appendix F.
+Universal Attack: (Right) All models share some percentage of universal attacks.
+Increasing the number and diversity among models could enhance their robustness against such attacks.
+RobustMath Ours ($M3$)
+Model OA AA ASR OA AA ASR $\Delta$ ASR
+Mistral 7B 10.3 18.7 0.0 33.0 0.0 100.0 +100.0
+MetaMath 7B 91.1 79.3 13.0 97.5 51.0 48.5 +35.5
+Llama-2 13B 2.3 8.3 0.0 11.0 0.0 100.0 +100.0
+WizardMath 13B 71.0 70.3 1.0 89.0 15.5 82.6 +81.6
+Vicuna 13B 46.3 51.7 0.0 68.0 2.0 97.5 +97.5
+CodeLlama 34B 31.3 10.3 67.1 8.5 0.0 100.0 +32.9
+MetaMath 70B 93.0 82.7 11.1 98.5 68.0 31.1 +20.0
+GPT-3.5 91.1 75.7 16.9 94.0 63.0 33.3 +16.4
+Average 54.6 49.6 13.6 62.4 24.9 74.1 +60.5
+Table 3: Comparison Result We calculate the average of each metric from both datasets and compare the result with rephrasing attack, RobustMath.
+Our method significantly outperforms the baseline in every model, with an average improvement of 60.5 ASR points.
+Universal Attacks
+Following the previous adversarial attack literature, we investigate whether there are adversarial examples exist in all LLMs, also known as universal attacks (Zou et al., 2023; Moosavi-Dezfooli et al., 2017).
+We count the number of adversarial examples from each model and calculated the percentage of common ones among them.
+We report the results on the right of Table 2 and observe a clear pattern of decreasing universal attacks with an increased number of models.
+This suggests that while all models are vulnerable to some degree of universal attacks, the percentage of such attacks decreases as more and diverse models are considered.
+We also observe that universal attacks varies significantly across different $M$s, with $M3$ having the lowest number of universal attacks.
+Efficient Attacks
+Scaling up the number of attacks results in a steady performance drop for all models.
+However, this scaling might not be viable for API-based models such as GPT-4 due to request limits and high costs.
 To circumvent this, we propose an efficient attack method for API-based models by using adversarial examples in a targeted manner.
-The idea is simple: we attack model $A$ (target model, GPT-4 in our case), with successful attacks from a cheaper model $B$ (e.g., open-source, lower API cost).
-We first select 50 problems along with their 100 adversarial examples running against GPT-4.
-Next, we select 50 successful attacks from a cheaper model and run them against GPT-4.
+The idea is simple: we attack model $A$ (target model, GPT-4 in our case), with adversarial examples from a cheaper model $B$ (e.g., open-source, lower API cost).
+We select 50 problems along with their 100 variations and run them against GPT-4, comparing the results with attacking GPT-4 using 50 adversarial examples from a cheaper model.
 Note that for some models, there are no correct responses, even with $M3$.
-Therefore, we only compare $M3$ with the models that has correct answers.
+Therefore, we only compare $M3$ with the models that have correct answers.
 We compare the results in Table 4 and observe a significant request reduction while achieving similar ASR.
-Method Req. RR. (%) ASR (%)
-Original 5000 0 10
-MetaMath 7B 1389 72 8
-WizardMath 13B 1745 65 6
-MetaMath 70B 672 87 8
-GPT-3.5-turbo 456 91 10
-Table 4: Efficient Attack: a targeted approach to attack high-cost models.
-We leverage successful attacks from cheaper models to attack GPT-4, achieving the same ASR with about 90% reduction in requests.
-4.7 Human Evaluation
-We conduct a human evaluation on the generated adversarial examples to ensure they are indeed valid.
-We randomly select one adversarial example from 20 problems from all three generation methods (totaling 60 examples) that GPT-3.5 failed on for GSM8K and presented them to three human evaluators to assess (i) correctness: whether the answer to the modified problem is correct; (ii) coherence: whether the problem is contextually coherent; and (iii) similarity: whether the newly generated problem’s difficulty level is similar to the original one.
-The annotators were asked to make a binary decision to determine whether the criteria described above were met.
-We present the average of their scores in Figure 4.
-Overall, we found that all human evaluators agreed with the correctness.
-$M3$ scores highly across all metrics, indicating that it indeed preserves the original problem’s coherence and difficulty level.
-On the other hand, the scores for coherence and similarity varies between generation methods, suggesting that while they all successfully perturbed the model, some of the examples may not be useful in an educational context.
-Figure 4: Human Evaluation: we report the average score among three annotators on the successful attacks from GPT-3.5-turbo.
-$M3$ shows the highest score across all metrics, indicating our best generation method correctly generate contextual coherent problems that preserve the original level of difficulty.
+Models Request Call Cost (\$) ASR $\Delta$ Request (%)
+MetaMath 7B 1,389 10.7 8.0 72.2
+WizardMath 13B 1,745 8.8 6.0 65.1
+MetaMath 70B 672 4.3 8.0 86.6
+GPT-3.5 456 2.8 10.0 90.9
+Target: GPT-4 5,000 29.3 10 -
+Table 4: Efficient Attack A targeted approach to attack high-cost models.
+We leverage adversarial examples from cheaper models to attack GPT-4, achieving the same ASR while reducing up to 90% request calls.
+Figure 2: Human Evaluation (Left) We report the average score among three annotators on the adversarial examples from GPT-3.5.
+$M3$ has the highest score across all metrics, indicating our best generation method correctly generates contextually coherent problems that preserve original difficulty.
+Transferability (Right) We present the adversarial example transferability among all models by comparing each model against all other models.
+Compared to the math-tuned and production models, the weaker models such as Mistral 7B and LLama2 13B exhibit significant vulnerability and a strong correlation among them.
 5 Analysis and Discussion
-In this section we conduct several analyses to better understand our generation methods and LLMs mathematical capabilities.
-5.1 Transferability
+In this section, we conduct analyses and discussions to better understand our generation methods and the mathematical capabilities of LLMs.
+Human Evaluation
+We conduct a human evaluation to ensure the our generation are indeed valid.
+Three human evaluators are asked to assess (i) correctness: whether the answer to the modified problem is correct; (ii) coherence: whether the problem is contextually coherent; and (iii) similarity: whether the newly generated problem’s difficulty level is similar to the original one.
+We randomly select one adversarial example from 20 problems that GPT-3.5 failed on for GSM8K.
+The annotators made binary decisions to determine whether the criteria were met.
+We present the average of their scores on the left of Figure 2.
+Overall, all human evaluators agree on the correctness.
+$M3$ scores highly across all metrics, indicating that it preserves the original problem’s coherence and difficulty level.
+Coherence and similarity are relatively low for $M2$ and $M3$, suggesting that while they might perturb the model, such generation may not be useful for the educational context.
+Transferability
 We investigate the transferability of adversarial examples among the models.
-We attack model $A$ with the successful attacks from model $B$ and calculate the number of common successful attacks between these two models.
-We present the GSM8K $M3$ result in Figure 5.
-Interestingly, we observe that Vicuna 13B, Llama2 13B, CodeLlama 34B, and Mistral 7B exhibit a high percentage of transferability, reaching 100%, indicating significant vulnerability and a strong correlation among these models.
-On the other hand, comparisons with MetaMath 70B and GPT-3.5 tend to have a lower transferability rate.
+We attack model $A$ with the adversarial examples from model $B$ and calculate the number of common adversarial examples between these two models.
+We present $M3$ result in the left of Figure 2.
+Interestingly, we observe that weaker models exhibit a high percentage of transferability, indicating a strong vulnerability correlation among these models.
+On the other hand, comparisons with strong models such as MetaMath 70B and GPT-3.5, tend to have a lower transferability rate.
 While it might suggest a form of resistance to adversarial examples that affect other models, it is not a measure of robustness as it could be that those models fail on entirely different examples.
 This finding suggests that certain LLMs might struggle with particular numbers or patterns.
-To further understand these phenomena, we conduct further analysis in §5.2.
-Figure 5: Transferability: we present the adversarial example transferability amount all models by comparing each model against all other models.
-Compare to the math-tuned and production models, the weaker models such as Mistral 7B and LLama2 13B exhibit significant vulnerability and a strong correlation amount them.
-5.2 Question Feature Analysis
-As a step towards understanding problems that are hard for LLMs, we conduct regression analysis to understand the effects of various features of the problems.
-More specifically, we construct features from the 20,000 MWPs using the generation method $M3$ (200 problems, each with 100 modifications).
-The input features include operation counts, value ranges, and node counts in the AST, resulting in a total of 51 input features.
-The features are plotted against the correctness of the model, with correct being $1$ and wrong being 00.
-We present the coefficients of the eight models with some selected features in Table 7 in the Appendix.
-The coefficients are only compared within each model, but not across models.
-A more suited metric to compare between models is accuracy, such as Table 1 and Figure 3.
-We present a set of graphs showing models’ accuracy versus some prominent features in Figure 7.
-Overall, we found that the value of the answer has a stronger correlation to the correctness.
-Somewhat surprisingly, most models have positive correlations between the number of constants and variables in a question and the correctness.
-The reason might be that the problems with more numbers have more straightforward operations (additions and subtractions).
-Additionally, we found that most high-performing models exhibit a negative correlation with the number of divisions in the problem, except for MetaMath 70B, and operation counts negatively correlate with correctness for all models.
-5.3 Code Analysis
-While the majority of generated Python code adheres to correct problem-solving logic and produces answers correctly, to ensure that the generated code is correctly being converted into ASTs, we manually validate the generated ASTs and identify the following errors: (1) Incorrect answers or code; (2) Use of loops, conditional, and comparison statements, and functions; (3) Use of expressions such as x = y // z + (y % z > 0); (4) Number misalignment, where the same number appears multiple times in the code or the problem.
-For GSM8K, 7 out of 200 problems were filtered out.
-For MultiArith, 15 out of 200 problems were filtered out.
-We discuss the methods of identifying and handling such errors in Appendix F.
-6 Conclusion
-In conclusion, we introduce a novel approach to generate MWPs that challenge the LLM’s math-solving abilities, thereby addressing the urgent need for methods that ensure fair educational evaluations.
-By utilizing AST to create adversarial examples, our generation method significantly degrades both open- and closed-source LLMs without compromising the problems’ original difficulty and coherence.
-We investigate the shared mathematical vulnerabilities among LLMs and propose a cost-effective method to attack high-cost API-based models, achieving up to a 90% decrease in request rates without compromising performance.
-Our automatic analyses shed light on the specific mathematical aspects that most influence LLM performance.
-This work not only paves the way for further research into developing more robust and ethically sound educational tools but also offers valuable insights into LLMs’ limitations, contributing to the ongoing discourse on the ethical use of LLMs in education.
-7 Limitations
-Perceived Difficulty Level: In this work, the correlation between the complexity of problems generated to challenge LLMs and the actual difficulty perceived by human students has not been empirically validated.
-It is plausible that modifications considered neutral or simple for LLMs could be perceived as more challenging by human students.
-We plan to further investigate this in future work.
-Imperfect Logical Coherence: While $M3$ produces similar numbers to the original problems, logical coherence is not perfect in few cases.
-For example, the phrases like “25 hour in a day” and “37 days in a month”.
-While this issue can be easily addressed by employing heuristic rules to associate certain numbers (e.g., 7, 24, 365) with specific entities (e.g., “days”, “hours”, “year”), we try to keep everything as generalized as possible without overfitting to specific cases.
-Cannot Alter Words: In this work, we only consider actual numeric values in a question; however, there are problems where words also represent numeric values.
-For example, a question from GSM8K is “A frog lays her eggs over a series of 4 days.
-The first day she lays 50 eggs.
-The second day, she doubles her production of eggs.
-The third day she lays 20 more than the second day, and the last day she doubles the first three days total.
-How many eggs did the frog lay over the span of the 4 days?” contains words like “doubles” and “three” that also represent numeric values.
-In the future, we plan to add a preprocessing step for the problems to extract as many numeric values from a question as possible, resulting in more diverse generations of problems.
+To further understand these phenomena, we conduct regression analysis as shown below.
+Question Feature Analysis
+To further understand the limitations of LLMs, we conduct a regression analysis to investigate the effects of various features of MWPs on the correctness of the model’s predictions.
+We construct a total of 51 input features from 20,000 $M3$ generated problems, including features such as operation counts, value ranges, and node counts in the AST.
+The features are plotted against the correctness of the model, with correct being 1 and wrong being 0.
+We present the coefficients of the all LLMs with some selected features in Table 7 in the Appendix.
+We observe that the most positively and negatively correlated features vary across different models, suggesting that our adversarial examples exploit different weaknesses of each model.
+For example, models exhibit varying performance with different answer value ranges.
+Mistral 7B and WizardMath 13B perform relatively well on problems with smaller answer values (e.g., in the range of [2, 8) and [8, 32)), while MetaMath 7B and Vicuna 13B show better performance on problems with answer values in the range of [32, 128).
+This suggests that different models may have learned to specialize in problems with specific range, possibly due to the varying distributions of the pre-training data.
+Overall, problems involving division and more operations tend to be more challenging for most LLMs.
+This confirms the intuition that complex problems requiring more mathematical operations are generally more difficult for LLMs to solve (Imani et al., 2023b).
+Visualizations are available in Figure 7 in the Appendix.
+Python Code Analysis
+While the majority of generated Python code adheres to correct problem-solving logic and produces answers correctly, to ensure that the generated code is correctly being converted into ASTs, we manually validate the generated ASTs and identify the following four types of possible errors: (1) Incorrect answers or code; (2) Use of loops, conditional, and comparison statements, and functions; (3) Use of expressions such as x = y // z + (y % z > 0); (4) Number misalignment, where the same number appears multiple times in the code or the problem.
+Encouragingly, for GSM8K and MultiArith, 96.5% and 92.5% of ASTs, respectively, are found to be valid and free from these errors.
+We discuss the methods of identifying and handling such errors in Appendix I.
+Why LLMs suffer From Such Simple Attack?
+While LLMs can be trained on vast amounts of math data, they may not have an inherent grasp of the underlying mathematical concepts and reasoning steps (Saxton et al., 2019; Lample & Charton, 2019).
+Instead, they likely rely on pattern and statistical associations learned from the training data to make predictions (Bender et al., 2021), which is also know as memorization (Carlini et al., 2022).
+Modifying the numbers in a problem may disrupt these learned patterns, causing the models to make errors.
+This is evidenced by the significant performance drops observed in models like WizardMath (89% to 20%) and Vicuna (76% to 4%) when presented with adversarial examples.
+Furthermore, test data contamination could also play a role - if the training corpus contains the test data, the models may be able to answer the original questions correctly but struggle to generalize to even subtle variations (Balloccu et al., 2024).
+6 Conclusion and Future work
+We introduce a novel approach to generate MWPs that challenge the LLM’s math-solving abilities, thereby addressing the urgent need for methods that ensure fair educational evaluations.
+By utilizing AST to create adversarial examples, we preserve problems’ original difficulty and coherence while significantly degrading both open- and closed-source LLMs and surpassing the previous rephrasing attack by an average of 60% ASR.
+We conduct a human evaluation to validate the adversarial examples and investigate universal attacks and attack transferability, proposing a cost-effective method to attack high-cost API-based models, achieving up to a 90% request reduction while maintaining the same performance.
+Our automatic regression analyses demonstrate that our adversarial examples exploit different weaknesses of each model, shedding light on the specific mathematical aspects that most influence individual LLM performance.
+This work not only paves the way for further investigation into the mathematical limitations among LLMs but also contributes to developing more robust and ethically sound educational tools to ensure fair evaluation, contributing to the ongoing discourse on the ethical use of LLMs in education.
+In this work, the correlation between the complexity that challenges LLMs and the actual difficulty perceived by humans has not been empirically validated.
+In future work, we aim to explore this and expand the methods to handle more complex problems.
 7 Ethics Statement
@@
 This approach risks amplifying the disparities between institutions or individuals with differing levels of technological access.
-9 Acknowledgements
-We are very grateful to Sandy Yeung and Jack Goffinet for the helpful initial project discussion.
-Thank you to the anonymous reviewers for their feedback.
-The research reported here was supported by the Learning Engineering Virtual Institute funded by leading education philanthropists and organizations through Grant G-23-2137070, to the University of Florida and its partner institutions.
-The opinions expressed are those of the authors and do not represent the views of the University of Florida, the partner institutions, or those of the philanthropists and organizations.
-Appendix A Prompt Template
+Appendix A Prompt Templates
 A.1 Code Generation Prompt
-[…]
+Code Generation Prompt Write a Python script that contains a step-by-step solution for a given math problem and its answer.
+Start the script by defining variables corresponding to the quantities mentioned in the problem.
+Print the final answer at the end of the script.
+Do not use if-else statements, floor divisions, and loops to solve the problem.
+Given problem: {problem} Answer: {answer} Python script:
 A.2 Zero-shot CoT
-[…]
+Evluation Prompt Solve a math problem. The solution ends with “the answer is (a number)” like “the answer is 1”. Question: {problem} Answer: Let’s think step by step.
 Appendix B Rephrasing Example
@@
 We associate the number in the original question with the variable node so that a change in the variable node will also change the number in the question.
-If a number appears multiple times in the code and the problem, then we associate them based on the specific procedures described in Section F.4.
+If a number appears multiple times in the code and the problem, then we associate them based on the specific procedures described in Section I.4.
 Appendix D Generation Details
@@
 Output: $v^{\prime}_{i}$
+Let $a_{i}\times 10^{b}_{i}\approx h=val(v_{i})$, where $a_{i}$ is an integer that is not divisible by 10 and has at most $c$ scientific digits.
 […]
@@
 […]
-                     Let $d$ be the number of digits $val(v_{i})$ has. if $d=1$ then
+                     Let $d$ be the number of digits $val(v_{i})$ has.
+                     if $d=1$ then
                           $a^{\prime}_{i}\sim\text{Uniform}(1,99)$
@@
                                               $a^{\prime}_{i}\sim\text{Pois}(a_{i})$
-$val(v^{\prime}_{i})\leftarrow a^{\prime}_{i}\times 10^{b}$ return $v^{\prime}_{i}$
+                                              $val(v^{\prime}_{i})\leftarrow a^{\prime}_{i}\times 10^{b}$
+return $v^{\prime}_{i}$
 Algorithm 1 Number Generation Based on Method
@@
 Output: $W^{\prime}=\{V^{\prime}_{1},\dots,V^{\prime}_{m}\}$, where $V^{\prime}_{j}=\{v^{\prime}_{1},\dots,v^{\prime}_{n}\}$
-$attempts\leftarrow 0,j\leftarrow 1$ $W^{\prime}=\emptyset$ while $attempts<N$ do
-      $V^{\prime}_{j}=\emptyset$ for $v_{i}\in V$ do
-           Obtain $v^{\prime}_{i}$ from Algorithm 1 with input $v_{i}$, $M$, and $c$ $V^{\prime}_{j}\leftarrow V^{\prime}_{j}\cup\{v^{\prime}_{i}\}$
-           $attempts\leftarrow attempts+1$ $S^{\prime}\leftarrow S$ Apply the new sequence of values from the variable nodes $V^{\prime}_{j}$ to $S^{\prime}$ $accept\leftarrow True$ for each node $s^{\prime}_{k}$ in $S^{\prime}$ do
+$attempts\leftarrow 0,j\leftarrow 1$
+$W^{\prime}=\emptyset$
+while $attempts<N$ do
+      $V^{\prime}_{j}=\emptyset$
+      for $v_{i}\in V$ do
+           Obtain $v^{\prime}_{i}$ from Algorithm 1 with input $v_{i}$, $M$, and $c$
+           $V^{\prime}_{j}\leftarrow V^{\prime}_{j}\cup\{v^{\prime}_{i}\}$
+           $attempts\leftarrow attempts+1$
+           $S^{\prime}\leftarrow S$
+           Apply the new sequence of values from the variable nodes $V^{\prime}_{j}$ to $S^{\prime}$
+           $accept\leftarrow True$
+           for each node $s^{\prime}_{k}$ in $S^{\prime}$ do
 […]
@@
                      if accept then
-                          $W^{\prime}\leftarrow W^{\prime}\cup\{V^{\prime}_{j}\}$ $j\leftarrow j+1$
+                          $W^{\prime}\leftarrow W^{\prime}\cup\{V^{\prime}_{j}\}$
+                          $j\leftarrow j+1$
                           return $W^{\prime}$
@@
 Method Question Answer
-Original
-Sarah makes 5 times more money per hour than
-Connor does. If Connor earns \$7.2 per hour, how
-much does Sarah make in an 8-hour day?
-288
-M3
-Sarah makes 7 times more money per hour than
-Connor does. If Connor earns \$8.0 per hour, how
-much does Sarah make in an 7-hour day?
-392
-M2
-Sarah makes 20 times more money per hour than
-Connor does. If Connor earns \$7.7 per hour, how
-much does Sarah make in an 8-hour day?
-1232
-M1
-Sarah makes 76780 times more money per hour than
-Connor does. If Connor earns \$7655.5 per hour, how
-much does Sarah make in an 12568-hour day?
-7387335796720
 Original
@@
 200411
-Table 5: Three examples of questions generated by the three generation methods.
+Table 5: Examples of questions generated by the three generation methods.
 Appendix E Rephrasing Attack Performance
@@
 MetaMath 70B 93.0 82.7 11.1
-GPT-3.5-Turbo 91.1 75.7 16.9
+GPT-3.5 91.1 75.7 16.9
 Table 6: Rephrasing Attack: As a baseline, we present the performance of RobustMath Zhou et al.
@@
 This suggests that rephrasing the MWPs may not generalize effectively across different LLMs.
-Appendix F Code Analysis
-F.1 Incorrect Answer from the Dataset
+Appendix F Incorrect Variant Count Distribution
+Figure 4: We present the distribution of the number of incorrect adversarial examples in different buckets.
+For strong models (e.g., GPT-3.5-Turbo, MetaMath 70B), around half of the problems have zero incorrect adversarial examples, while for weak models (e.g., Llama-2 13B, CodeLlama 34B), most problems has more than 90 incorrect adversarial examples.
+Appendix G Number of Generation
+Figure 5: Possible Generations: with 30,000 attempts for each given constraint, we calculate the average number of adversarial examples generated for each problem.
+Appendix H Number of Attacks
+Figure 6: We report the model performance given various number of attacks. For several models, only 10 attacks are enough to degrade the performance.
+Appendix I Code Analysis
+I.1 Incorrect Answer from the Dataset
 There are instances where the dataset-provided answers might be wrong.
-An example of such an instance is the following question from the MultiArith dataset: "Paige’s team won their dodgeball game and scored 41 points total.
-If Paige scored 11 of the points and everyone else scored 6 points each, how many players were on her team?", which the dataset provided 5 as the answer.
+An example of such an instance is the following question from the MultiArith dataset: ”Paige’s team won their dodgeball game and scored 41 points total.
+If Paige scored 11 of the points and everyone else scored 6 points each, how many players were on her team?”, which the dataset provided 5 as the answer.
 However, the correct answer should be 6, as Paige is on the team as well.
@@
 To automatically detect such instances, the value of the final answer node will be compared to the answer given by the dataset, and the question will be filtered out for further generations of problems if the two answers don’t match.
-F.2 Incorrect Answer from GPT-4
+I.2 Incorrect Answer from GPT-4
 Similarly, there are also instances where GPT-4 generated codes that provided the wrong answer to the given question.
-One such question is from the GSM8K dataset: "Jen decides to travel to 3 different countries.
+One such question is from the GSM8K dataset: ”Jen decides to travel to 3 different countries.
 He has to pay \$400 for the supplies he needs, in total.
 The tickets for travel cost, in total, 50% more than the supplies.
-How much does travel cost?" The correct answer should be $400+400\times 1.5=1000$, but the code generated by GPT-4 outputted \$600.
+How much does travel cost?” The correct answer should be $400+400\times 1.5=1000$, but the code generated by GPT-4 outputted \$600.
 Similar to the previous error, instances where the code outputs the wrong answers will be detected by comparing them to answers given by the dataset, and the instances will be discarded from generating math problems.
-F.3 Contains Control Flow Statements
-Although specified in the prompt, GPT-4 might generate codes that include control flow statements (including "if", "for", and "while" statements).
+I.3 Contains Control Flow Statements
+Although specified in the prompt, GPT-4 might generate codes that include control flow statements (including ”if”, ”for”, and ”while” statements).
 Codes that contain control flow statements are considered errors and discarded because the codes are usually only applicable to the original instance of the question.
 Furthermore, the code might not halt for some combinations of numbers, resulting in no answers associated with the generated instances.
-F.4 Number Misalignment
+I.4 Number Misalignment
 To ensure the correct generation of new questions, numbers in the original question need to be associated with numbers in the code.
@@
 Then, random numbers are generated to combine the equivalences of correspondences.
-For example, given that the question is "Mary has 5 apples and 5 oranges, how many fruits does Mary have?".
-Although the two "5"s correspond to different representations in the question, due to the communicative nature of addition, the two "5"s are interchangeable in the syntax tree.
+For example, given that the question is ”Mary has 5 apples and 5 oranges, how many fruits does Mary have?”.
+Although the two ”5”s correspond to different representations in the question, due to the communicative nature of addition, the two ”5”s are interchangeable in the syntax tree.
 Thus, the two pairs of correspondence for this question can be combined, and there is no ambiguity in the grounding.
@@
 The correspondence that has the same answer will be used, and if no correspondence has the correct answer, the question will be discarded.
-An example of this type of question is: "The pizzeria sells small pizzas for \$2 and large pizzas for \$8.
+An example of this type of question is: ”The pizzeria sells small pizzas for \$2 and large pizzas for \$8.
 They sold \$40 in pizzas.
-If they sold 8 small pizzas, how many large pizzas did they sell?".
+If they sold 8 small pizzas, how many large pizzas did they sell?”.
 One valid and distinguishing sequence of numbers for the question is $(2,3,31,5)$, which will result in different answers $5$ and $7$.
@@
     There is a number $c_{i}$, such that $C_{q}(c_{i})>C_{c}(c_{i})$. Similar to the reasoning for the first case, these kinds of questions are eliminated for problem generations.
-Appendix G Feature Analysis
-Features MetaMath-7B-V1.0 vicuna-13b-v1.5 CodeLlama-34b-hf gpt-3.5-turbo-1106
+Appendix J Feature Analysis
+Features MetaMath 7B Vicuna 13b CodeLlama 34b GPT-3.5
 Addition Count -0.0032 0.0080 -0.0146 0.0098
@@
 Constant 0.2840 0.1840 0.0328 0.3919
-Features Llama-2-13b-hf MetaMath-70B-V1.0 Mistral-7B-v0.1 WizardMath-13B-V1.0
+Features Llama 2 13b MetaMath 70B Mistral 7B WizardMath 13B
 Addition Count -0.0163 -0.0287 0.0102 -0.0118
@@
 Constant 0.0205 0.3100 0.0805 0.2800
-Table 7: Coefficients for some selected features after fitting linear regression models to the features of problems and the correctness of the eight models.
-Note that coefficients should not be compared across models, only with other features within the same model.
-For example, we can observe that MetaMath-7B performs well when the answer is in the range [2, 512), while performance declines when the answer is in the range [512, 8192).
+Table 7: Coefficients of regression analysis on selected features for each model.
+Positive coefficients indicate that the model performs better on problems with the corresponding feature, while negative coefficients indicate the opposite.
+The most positive and negative correlation coefficients for each model are bolded, and the second most positive and negative correlation coefficients are underlined.
+We observe that the most positively and negatively correlated features vary across different models, suggesting that our adversarial examples exploit different weaknesses of each model.
+For example, Mistral 7B and WizardMath 13B perform relatively well on problems with smaller answer values (e.g., in the range of [2, 8) and [8, 32)), while MetaMath 7B and Vicuna 13B show better performance on problems with answer values in the range of [32, 128).
 The observation matches the accuracy shown in Figure 7.
-Figure 7: The accuracy of eight models is plotted against four relevant features when tasked with all questions generated by method M3.
-The top left graph shows the accuracy of models against problems with various counts of multiplications, and the top right graph compares problems with various counts of all operations, including binary and unary.
-The middle graph shows the accuracy of models with different ranges of generated numbers, and the bottom graph shows the accuracy with different ranges of final answers.
+Note that coefficients should not be compared across LLMs, only with other features within the same model.
+A more suited metric to compare between LLMs is accuracy, such as Table 1 and Figure 6.
+Figure 7: The accuracy of LLMs is plotted against five relevant features when tasked with all questions generated by $M3$.
+The top left graph shows the accuracy of models against problems with various counts of multiplications, and the top right graph compares problems with various counts of all operations, including both binary and unary operations.
+The middle left graph shows the accuracy against token counts of the final answer, and the middle right graph shows the accuracy against the count of nodes in the generated AST.
+The bottom two graphs shows the accuracy of models with different ranges of generated numbers and the range of final answers.

### v3

@@
-LLM-Resistant Math Word Problem Generation via Adversarial Attacks
+Adversarial Math Word Problem Generation
 Abstract
@@
 We identify shared vulnerabilities among LLMs and propose a cost-effective approach to attack high-cost models.
-Additionally, we conduct automatic analysis on math problems and investigate the cause of failure, offering a nuanced view into model’s limitation.
+Additionally, we conduct automatic analysis to investigate the cause of failure, providing further insights into the limitations of LLMs.
+[^1] Data and code are available: https://github.com/ruoyuxie/adversarial_mwps_generation
+Adversarial Math Word Problem Generation
+Roy Xie      Chengxuan Huang      Junlin Wang      Bhuwan Dhingra Duke University {ruoyu.xie, jonathan.huang, junlin.wang2}@duke.edu {bdhingra}@cs.duke.edu
 1 Introduction
@@
 Our aim is not just to challenge LLMs but to do so in a way that reflects real-world educational standards to ensure the MWPs remain educationally valuable and relevant.
-Figure 1: Method Overview Given a MWP that an LLM can correctly solve, our method first transforms it into Python code.
-The Python code then is converted into an AST representation which is used to generate adversarial problems by modifying the numeric values in the problem.
+Figure 1: Method Overview: Given a MWP that an LLM can correctly solve, our method first transforms it into Python code.
+The Python code then is converted into an AST representation, which is used to generate adversarial problems by modifying the numeric values in a controllable manner.
 We place constraints on the nodes of the AST to ensure that the modified problem maintains the same difficulty level as the original problem.
@@
 We then define educational constraints for each node to ensure all the desired properties are preserved for the generated new problem.
-In this work, we evaluate several major LLMs and demonstrate the effectiveness of our method by achieving a significant attack success rate (ASR).
-Our approach outperforms the previous rephrasing attack by an average of 60% ASR.
-We investigate universal attacks and attack transferability, proposing a cost-effective approach to attack high-cost models (e.g., GPT-4) by reducing API request calls by 90% without sacrificing performance.
+In this work, we evaluate several LLMs and demonstrate the effectiveness of our method by achieving a significant attack success rate (ASR).
+Our approach outperforms the previous rephrasing attack by an average of 62% ASR.
+We investigate universal attacks and attack transferability, proposing a cost-effective approach to attack high-cost models (e.g., GPT-4) by reducing API request calls by 90% while achieving high performance.
 We conduct human evaluations to verify that our generated problems indeed preserve the original coherence and difficulty.
-Furthermore, we perform a regression analysis and find that our adversarial examples exploit different weaknesses of each model, offering valuable insights and a nuanced view into model’s limitation.
+Furthermore, we perform a regression analysis and find that our adversarial examples exploit different weaknesses of each model, offering valuable insights into LLM’s limitation.
 2 Background and Related Work
@@
 LLMs Math-solving Ability
-Our work also relates closely to LLMs math reasoning ability (OpenAI, 2023; Yu et al., 2023; Xu et al., 2023).
+Our work also closely relates to LLM’s math reasoning ability (Yu et al., 2023; Xu et al., 2023).
 Studies found that LLMs can significantly improve their math-solving ability through prompt engineering (Yu et al., 2023; Imani et al., 2023a), such as chain-of-thought (CoT) (Wei et al., 2022).
-Converting the MWPs’ solving steps into code or symbolic representations can also improves LLM’s performance (Li et al., 2023; He-Yueya et al., 2023; Gao et al., 2023).
+Converting the MWPs’ solving steps into symbolic representations can also improves LLM’s performance (Li et al., 2023; He-Yueya et al., 2023; Gao et al., 2023).
 [^2] While different prompting strategies exist, our focus is to generate adversarial examples that LLMs are unable to solve.
 Therefore, we try to minimize prompt engineering by using a unified prompt with zero-shot CoT for all evaluations.
-Prompt can be found in Appendix A.2.
+The prompt can be found in the appendix A.2.
 Adversarial Attacks on MWPs
 Adversarial attacks on LLMs involve modifying prompts to alter their behavior (Zhang et al., 2020; Zou et al., 2023; Carlini et al., 2023).
-Similarly, we modify MWPs to cause LLMs to output incorrect answers.
+In this work, we modify MWPs to cause LLMs to output incorrect answers.
 Bubeck et al.
-(2023) conducted limited memorization tests on MWPs by randomly changing numeric values between -10 and 10, suggesting that state-of-the-art LLMs do not solely rely on memorization but apply general solution methods.
+(2023) conducted limited memorization tests on MWPs by randomly changing numeric values, suggesting that state-of-the-art LLMs do not solely rely on memorization but apply general solution methods.
 However, their study’s sample size was relatively small, and we demonstrate that modifying numbers in MWPs causes LLMs to fail on a larger scale.
@@
 (2023) rephrased MWPs by changing words while preserving numeric values.
-This approach risks altering the original context and introducing inconsistencies and problematic content (see Appendix B), necessitating extensive human validation.
-Therefore, we focus on changing numerical values, preserving the underlying logic.
-To the best of our knowledge, this is the first study to undertake such an endeavor.
+This approach risks altering the original context and introducing inconsistencies and problematic content (see Appendix B), requiring extensive human validation.
+Therefore, we focus on altering the numerical values in this work and preserve the underlying logic.
 3 Methodology
-MWPs are presented using natural language, which creates challenges for systematic structural and syntactic modifications.
+MWPs are presented in natural language, which creates challenges for systematic structural and syntactic modifications.
 In this section, we describe our approach to generating adversarial MWPs.
@@
 We denote an original problem-answer pair as $p=(x,y)$, where $x$ is a sequence of tokens $x_{1},\dots,x_{n}$ in the problem and $y$ is its ground-truth answer.
-We define $G$ as a ground-truth function that computes the correct answer for a math problem [^2] In our case $G$ is generated Python code.
-and $F$ as a function which maps the elements of a sequence as: $F(\tilde{x}_{i})\sim\mathbb{R}$ (i.e., a random real number) if $x_{i}$ is numeric and $F(\tilde{x_{i}})=x_{i}$ otherwise.
+We define $G$ as a ground-truth function that computes the correct answer for a math problem and $F$ as a function which maps the elements of a sequence as: $F(\tilde{x}_{i})\sim\mathbb{R}$ (i.e., a random real number) if $x_{i}$ is numeric and $F(\tilde{x_{i}})=x_{i}$ otherwise.
+[^3] In our case $G$ is generated Python code.
 We denote the set of all possible adversarial modifications to it as:
@@
 To structurally modify MWPs, we first use GPT-4 to generate the Python code that reflects the solution steps given the problem and its final answer.
-[^3] We only require the problem and its final answer, which is almost always given in the context of math problems.
-Prompt can be found in Appendix A.1.
-As a sanity check, we select 200 MWPs each from GSM8K (Cobbe et al., 2021) and MultiArith (Roy & Roth, 2015) and check if the generated program results in correct answers.
-Given the execution steps in code are almost always deterministic, encouragingly, we found only 1% and 2.5% generated codes have incorrect logic or answers in GSM8K and MultiArith, respectively.
-We further investigate the generated Python code in §5.
-Next, we utilize AST, a tree for the abstract syntactic structure of programming language code, to convert the generated Python code into a tree representation.
+[^4] This is an one-time operation.
+We only require the problem and its final answer, which is almost always given in the context of math problems.
+The prompt is in the Appendix A.1.
+Next, we utilize the AST, a tree structure for programming language code, to convert the generated Python code into a tree representation for controllable new problem generation.
 We build the ASTs by traversing through the Python code and constructing nodes corresponding to each statement.
@@
 Each AST has mainly two types of nodes: operation nodes, which carry out the operations in the tree and are the non-leaf nodes, and variable nodes, which correspond to the numeric values from the problem and are the leaf nodes.
-We discuss the nodes in AST in detail in Appendix C.
+We discuss the nodes in ASTs in detail in Appendix C and conduct analysis on the quality of generated Python code and ASTs in §5.
 3.2 Adversarial Example Generation
-In an adversarial setting, attackers can perform multiple attacks on the system, and any successful attack signifies the possible flaws of the system (Carlini & Wagner, 2017).
+In an adversarial setting, attackers can perform multiple rounds of attack on the system, and any successful attack signifies the possible flaws of the system (Carlini and Wagner, 2017).
 Similarly, we generate adversarial examples which are different versions of the original MWPs to attack LLMs.
@@
 For example, changing a problem from “Jack ate 2 out of a total of 5 apples” to “Jack ate 5 out of a total of 2 apples” not only introduces the concept of negative numbers but also alters the problem’s logical structure and coherency.
-Furthermore, such modifications can lead to unrealistic scenarios that deviate from the intended learning objectives and may cause confusion.
+Such modifications can create unrealistic scenarios, potentially confusing students and leading to ineffective learning experiences Vilenius-Tuohimaa et al.
+(2008).
 Filtering Constraints
-To minimize the difference between the original MWP and its adversarial counterpart while adhering to the educational context, we define a set of Boolean constraints for each node to control the generation quality.
+To minimize the difference between the original MWP and its adversarial counterpart while adhering to the educational context, we define a set of Boolean constraints for each node in the AST to control the generation quality.
 Given the original node value $h$ and its new value $h^{\prime}$, a newly generated problem is valid if and only if $h^{\prime}$ satisfies the same set of constraints that $h$ has, for all nodes.
@@
 Depending on the numbers $h$, $a$, and $b$, the generation methods are defined as follows:
-  • $M1$: Free Generation: Allowing the generated problem to have a wide range of numbers with minimal constraints, with each $h^{\prime}$ to be $a^{\prime}\times 10^{b}$.
+  • $M1$ Free Generation: Allowing the generated problem to have a wide range of numbers with minimal constraints, with each $h^{\prime}$ to be $a^{\prime}\times 10^{b}$.
 Regardless of the value of $a$, $a^{\prime}$ is drawn from a uniform distribution between $1$ and $10^{c}$.
 If the variable node $v$ is a divisor of a division node, then $a^{\prime}$ is drawn from a uniform distribution between $1$ and $\left\lfloor 10^{c/2}\right\rfloor$.
-  • $M2$: Count of Digits: Constraining the generated value $h^{\prime}$ has the same number of digits as $h$.
+  • $M2$ Count of Digits: Constraining the generated value $h^{\prime}$ has the same number of digits as $h$.
 For example, for $h=100$, some possible $h^{\prime}$ are 942, 589, or 264.
@@
 For decimal variable values like $h=1.25$, we use $a=125$ to generate numbers like $a^{\prime}=473$, then convert them back to $h^{\prime}=4.73$.
-  • $M3$: Count of Scientific Numbers: Constraining the new value shares a similar scientific digit count and is within a similar range to the original value.
+  • $M3$ Count of Scientific Numbers: Constraining the new value shares a similar scientific digit count and is within a similar range to the original value.
 For example, for $h=1500$, $h^{\prime}\sim\text{Pois}(h)$ can be 1700, 800, 1200, where Pois is the Poisson distribution.
@@
 $M1(A(x,y))\supseteq M2(A(x,y))\supseteq M3(A(x,y)).$ (2)
-The more restrictive a method is, the closer the difficulty levels and coherence are between the an adversarial example and its original version.
-Therefore, we use $M3$ as our main generation method as it ensures the adversarial examples comply with all constraints to maintain the original difficulty and coherence.
-Although generations from $M1$ and $M2$ may not fit in the educational context, we are interested in studying LLMs’ math-solving abilities by simply changing the numeric values.
+The more restrictive a method is, the closer the difficulty levels and coherence remain between an adversarial example and its original version.
+Thus, we use $M3$ as our main generation method since it ensures that adversarial examples adhere to all constraints, maintaining original difficulty and coherence.
+While $M1$ and $M2$ generations may not fit the educational context, we aim to study LLMs’ math-solving abilities by simply altering numeric values.
 We discuss how different methods impact model performance in §4.2.
-We describe each generation method in more detail in Appendix D and present the generated examples based on three methods in Table 5.
+We present detailed descriptions of each generation method in Appendix D and their generated examples in Table 6.
 4 Experiments and Results
@@
 Datasets
-We generate problem variants from GSM8K (Cobbe et al., 2021) and MultiArith (Roy & Roth, 2015).
-Both datasets are commonly used for evaluating LLMs’ mathematical capabilities, and MultiArith is sourced from math worksheets used by elementary school students to practice math problems (Roy et al., 2015).
+We generate problem variants from GSM8K (Cobbe et al., 2021) and MultiArith (Roy and Roth, 2015).
+Both datasets are commonly used for evaluating LLMs’ mathematical capabilities, and MultiArith is sourced directly from math worksheets used by elementary school students to practice math problems (Roy et al., 2015).
 [^5] While MWPs exhibit different levels of difficulty, we focus on MWPs at this particular difficulty level as a proof of concept in this work.
-Models
-We conduct experiments on 7 open-source models: MetaMath 7B, 70B (Yu et al., 2023), Mistral 7B (Jiang et al., 2023), Llama 2 13B (Touvron et al., 2023), WizardMath 13B (Xu et al., 2023), Vicuna 13B (Chiang et al., 2023), CodeLlama (Roziere et al., 2023), and 2 closed-source models: GPT-4 (OpenAI, 2023) and GPT-3.5 (OpenAI, 2022).
-[^5] Specifically, we use GPT-4-0125-preview and GPT-3.5-turbo-0125.
-For simplicity, we refer to them as GPT-4 and GPT-3.5 for the rest of the paper.
-The model selection was largely based on the LLMs’ math-solving ability.
-Our intention is to evaluate a wide range of LLM performances, including the math-tuned LLMs such as MetaMath and WizardMath, popular open-source LLMs such as Llama 2 and Mistral, and the API-based GPT models.
-Metrics
-We follow the previous adversarial attack literature to measure the attacks, which looks for at least one perturbation that fools the model (Croce et al., 2020; Pruthi et al., 2019; Jia & Liang, 2017).
-A problem is considered incorrect if it has at least one incorrect variation.
-Such variation is also known as an adversarial example.
-Given a set of original problem-answer pairs $P$ and a LLM $L$, we define: [^6] OA, AA, ASR are reported in % in this work.
-  • Original Accuracy (OA): the accuracy of $L$ on the original problems,
-    $\text{OA}(L)=\frac{\sum_{(x,y)\in P}\mathbf{1}\{L(x)=y\}}{|P|}.$ (3)
-  • Attack Accuracy (AA): the accuracy of $L$ on the adversarial modifications,
-    $\text{AA}(L)=\frac{\sum_{(x,y)\in P}\mathrm{1}\left[\forall(\tilde{x},\tilde{y})\in A(x,y):L(\tilde{x})=\tilde{y}\right]}{|P|}.$ (4)
-  • Attack Success Rate (ASR): relative decrease in accuracy due to adversarial modifications,
-    $\text{ASR}(L)=\frac{\text{OA}(L)-\text{AA}(L)}{\text{OA}(L)}.$ (5)
 MultiArith GSM8K
@@
 Model OA AA ASR AA ASR AA ASR OA AA ASR AA ASR AA ASR
-Mistral 7B 37 0 100 0 100 0 100 29 0 100 0 100 0 100
-MetaMath 7B 100 74 26 10 90 0 100 95 28 71 9 91 0 100
-WizardMath 13B 89 20 78 5 94 0 100 89 11 88 2 98 0 100
-Vicuna 13B 76 4 95 1 99 0 100 60 0 100 0 100 0 100
-Llama 2 13B 12 0 100 0 100 0 100 10 0 100 0 100 0 100
-CodeLlama 34B 11 0 100 0 100 0 100 6 0 100 0 100 0 100
-MetaMath 70B 99 86 13 30 70 0 100 98 50 49 17 83 0 100
-GPT-3.5 97 74 24 47 52 0 100 91 52 43 31 66 0 100
-Average 65.1 32.3 66.9 11.6 88.1 0 100 59.8 17.6 81.4 7.4 92.3 0 100
-Table 1: Main Attacks We generate MWPs based on different generation methods.
-Overall, simply changing the numeric values in the problems causes a consistent performance drop across all LLMs, even with the most restrictive generation method.
+Mistral 7B 37.0 0.0 100.0 0.0 100.0 0.0 100.0 29.0 0.0 100.0 0.0 100.0 0.0 100.0
+MetaMath 7B 100.0 74.0 26.0 10.0 90.0 0.0 100.0 95.0 28.0 71.0 9.0 91.0 0.0 100.0
+Llama 3 8B 17.0 0.0 100.0 0.0 100.0 0.0 100.0 21.0 0.0 100.0 0.0 100.0 0.0 100.0
+Llama 2 13B 12.0 0.0 100.0 0.0 100.0 0.0 100.0 10.0 0.0 100.0 0.0 100.0 0.0 100.0
+WizardMath 13B 89.0 20.0 78.0 5.0 94.0 0.0 100.0 89.0 11.0 88.0 2.0 98.0 0.0 100.0
+Vicuna 13B 76.0 4.0 95.0 1.0 99.0 0.0 100.0 60.0 0.0 100.0 0.0 100.0 0.0 100.0
+CodeLlama 34B 11.0 0.0 100.0 0.0 100.0 0.0 100.0 6.0 0.0 100.0 0.0 100.0 0.0 100.0
+MetaMath 70B 99.0 86.0 13.0 30.0 70.0 0.0 100.0 98.0 50.0 49.0 17.0 83.0 0.0 100.0
+GPT-3.5 97.0 74.0 24.0 47.0 52.0 0.0 100.0 91.0 52.0 43.0 31.0 66.0 0.0 100.0
+Average 60.0 28.7 70.7 10.3 78.3 0.0 100.0 55.4 15.9 83.4 6.6 93.1 0.0 100.0
+Table 1: Main Attacks: Performance of three different generation methods.
+Simply changing numeric values consistently cause performance drop across all LLMs, even with the most restrictive generation method.
+Models
+We conduct experiments on the following open-source models: MetaMath 7B, 70B (Yu et al., 2023), Mistral 7B (Jiang et al., 2023), Llama 3 8B (AI@Meta, 2024), Llama 2 13B (Touvron et al., 2023), WizardMath 13B (Xu et al., 2023), Vicuna 13B (Chiang et al., 2023), and CodeLlama (Roziere et al., 2023).
+Additional, we evaluate two closed-source models: GPT-4 (OpenAI, 2023) and GPT-3.5 (OpenAI, 2022).
+[^6] GPT-4-0125-preview and GPT-3.5-turbo-0125.
+The model selection was largely based on the LLMs’ math-solving ability.
+Our intention is to evaluate a wide range of LLM performances, including the math-tuned LLMs such as MetaMath and WizardMath, popular open-source LLMs such as Llama 3 and Mistral, and the API-based GPT models.
+Metrics
+We follow the previous adversarial attack literature to measure the attacks, which looks for at least one perturbation that fools the model (Croce et al., 2020; Pruthi et al., 2019; Jia and Liang, 2017).
+A problem is considered incorrect if it has at least one incorrect variation, also known as an adversarial example.
+Given a set of original problem-answer pairs $P$ and a LLM $L$, we define: [^7] OA, AA, ASR are reported in % in this work.
+  • Original Accuracy (OA): the accuracy of $L$ on the original problems,
+    $\text{OA}(L)=\frac{\sum_{(x,y)\in P}\mathrm{1}\{L(x)=y\}}{|P|}.$ (3)
+  • Attack Accuracy (AA): given an indicator function $I_{xy}$:
+    $I_{xy}=\mathrm{1}\left[\forall(\tilde{x},\tilde{y})\in A(x,y):L(\tilde{x})=\tilde{y}\right],$ (4)
+    $L$’s accuracy on the adversarial examples:
+    $\text{AA}(L)=\frac{\sum_{(x,y)\in P}I_{xy}}{|P|}.$ (5)
+  • Attack Success Rate (ASR): relative decrease in accuracy due to adversarial modifications,
+    $\text{ASR}(L)=\frac{\text{OA}(L)-\text{AA}(L)}{\text{OA}(L)}.$ (6)
 Baseline
-As a baseline, we examine the effectiveness of a commonly used rephrasing approach, where the modifications are centered around the words within the MWPs.
+As a baseline, we evaluate a common rephrasing approach focusing on modifying words within MWPs.
 Zhou et al.
@@
 Problem Variation Generation
-We set the number of attempts to generate problem variants to be 30,000 and present the average number of generated variants per problem for each $M$ in Appendix G.
+We set the number of attempts to generate problem variants to be 30,000 and present the average number of generation per problem for each $M$ in Appendix G.
 While on average each problem generates over a thousand variants, it is worth noting that the actual number of possible generations varies based on the original values and the number of variable nodes in a problem.
 [^8] For example, a problem with only one digit value and one variable node has a potential maximum of 9 possible modifications under $M3$.
-Given this reason and cost considerations, in this work we equally select 100 random problems from MultiArith and GSM8K and generate 100 variants for each selected problem in each $M$.
-Despite the seemingly small number of selected problems, the total data points are up to 60,000.8883 generation methods $\times$ 200 selected problems $\times$ 100 variants = 60,000.
-Main Attacks
-We report the model performance from all $M$s in Table 1.999We do not run the full dataset against GPT-4 due to cost considerations.
-Instead, we propose a cost-effective approach to query expensive models in Efficient Attack section.
-Ranging from the most lenient method, $M1$, where the numbers could be fairly wild, to the most stringent one, $M3$, we observe a consistent accuracy drop across all models, even for the strong models (e.g., math-tuned, larger size, or API-based).
-For the most original-problem-like generation, $M3$, we observe that weak models (e.g., smaller size or general-purpose) fail to generate any correct answers in every case.
-The left of Table 2 shows each model’s average percentage of incorrect variants.
-It is worth noting that these percentages are highly correlated with the reported ASR, with GPT-3.5 and MetaMath being the most robust, and CodeLLama and LLama 2 being the least robust models.
-Furthermore, we compare $M3$ with rephrasing attacks in Table 3.
-Our method significantly outperforms the baseline in every model, resulting in 60.5 ASR point improvement on average.
-We also analyze how ASR changes given different numbers of attacks in Appendix H.
-Interestingly, for several models, just 10 attacks are enough to degrade the performance significantly.
+Given this reason and cost considerations, in this work we randomly select 100 random problems from MultiArith and GSM8K and generate 100 variants for each selected problem in each $M$.
+Despite the seemingly small number of selected problems, the total data points are up to 60,000.9993 generation methods $\times$ 200 selected problems $\times$ 100 variants = 60,000 data points.
 Models
-AE. (%)
+Avg. (%) $\downarrow$
+CodeLlama 34B 91.8
+Llama 2 13B 91.6
+Llama 3 8B 78.3
 Mistral 7B 70.5
+Vicuna 13B 50.3
+WizardMath 13B 21.0
 MetaMath 7B 15.2
-Llama 2 13B 91.6
-WizardMath 13B 21.0
-Vicuna 13B 50.3
-CodeLlama 34B 91.8
 MetaMath 70B 7.9
 GPT-3.5 6.9
-Model Ct. M3 (%) M2 (%) M1 (%)
-Mistral 7B 1 100.0 100.0 100.0
-MetaMath 7B 2 70.0 90.0 100.0
-Llama 2 13B 3 67.0 90.0 100.0
-WizardMath 13B 4 52.0 80.0 100.0
-Vicuna 13B 5 44.0 78.0 100.0
-CodeLlama 34B 6 21.0 67.0 99.0
-MetaMath 70B 7 9.0 49.0 87.0
-GPT-3.5 8 9.0 46.0 83.0
-Table 2: Incorrect Variants (Left) The average percentage of incorrect variants per problem for each model.
-A detailed count distributions can be found in Appendix F.
-Universal Attack: (Right) All models share some percentage of universal attacks.
-Increasing the number and diversity among models could enhance their robustness against such attacks.
+Table 2: Incorrect Variants: The average percentage of incorrect variants per problem for each model in $M3$. Appendix F shows detailed count distributions.
 RobustMath Ours ($M3$)
@@
 MetaMath 7B 91.1 79.3 13.0 97.5 51.0 48.5 +35.5
-Llama-2 13B 2.3 8.3 0.0 11.0 0.0 100.0 +100.0
+Llama 3 8B 22.0 30.0 0.0 19.0 0.0 100.0 +100.0
+Llama 2 13B 2.3 8.3 0.0 11.0 0.0 100.0 +100.0
 WizardMath 13B 71.0 70.3 1.0 89.0 15.5 82.6 +81.6
@@
 GPT-3.5 91.1 75.7 16.9 94.0 63.0 33.3 +16.4
-Average 54.6 49.6 13.6 62.4 24.9 74.1 +60.5
-Table 3: Comparison Result We calculate the average of each metric from both datasets and compare the result with rephrasing attack, RobustMath.
-Our method significantly outperforms the baseline in every model, with an average improvement of 60.5 ASR points.
+Average 51.0 47.4 12.1 57.6 22.2 77.0 +62.0
+Table 3: Comparison Result: We calculate the average of each metric from both datasets and compare the result with rephrasing attack, RobustMath.
+Our method significantly outperforms the baseline in every model, with an average improvement of 62 ASR points.
+Main Attacks
+We report the model performance from all $M$s in Table 1.101010We do not run the full dataset against GPT-4 due to cost considerations.
+Instead, we propose a cost-effective approach to query expensive models in Efficient Attack section.
+Ranging from the most lenient method, $M1$, where the numbers could be fairly wild, to the most stringent one, $M3$, we observe a consistent performance drop across all models, even strong ones (e.g., math-tuned, larger size, or API-based).
+For the most original-problem-like generation, $M3$, weak models (e.g., smaller size or general-purpose) fail to generate any correct answers.
+Table 2 shows the average percentage of incorrect variants in each model, which is highly correlated with the reported ASR, with GPT-3.5 and MetaMath being the most robust and CodeLLama and LLama 2 the least.
+Furthermore, comparing $M3$ with rephrasing attacks in the Table 3, our method significantly outperforms the baseline, resulting in a 62 ASR point improvement on average.
+We also analyze how ASR changes given different numbers of attacks in Appendix I.
+Interestingly, for several models, just 10 attacks are enough to significantly degrade performance.
+Model Ct. M3 (%) M2 (%) M1 (%)
+Mistral 7B 1 100.0 100.0 100.0
+MetaMath 7B 2 70.0 90.0 100.0
+Llama 3 8B 3 67.0 87.0 100.0
+Llama 2 13B 4 67.0 87.0 100.0
+WizardMath 13B 5 49.0 80.0 100.0
+Vicuna 13B 6 44.0 77.0 100.0
+CodeLlama 34B 7 19.0 65.0 99.0
+MetaMath 70B 8 9.0 49.0 87.0
+GPT-3.5 9 9.0 46.0 83.0
+Table 4: Universal Attack: Universal attacks are shared among a number of models.
+Increasing the count of models being considered decreases the universal attacks, with $M3$ consistently showing the lowest percentages.
 Universal Attacks
@@
 We count the number of adversarial examples from each model and calculated the percentage of common ones among them.
-We report the results on the right of Table 2 and observe a clear pattern of decreasing universal attacks with an increased number of models.
+We report the results on Table 4 and observe a clear pattern of decreasing universal attacks with an increased number of models.
 This suggests that while all models are vulnerable to some degree of universal attacks, the percentage of such attacks decreases as more and diverse models are considered.
 We also observe that universal attacks varies significantly across different $M$s, with $M3$ having the lowest number of universal attacks.
+Models Req Call Cost (\$) ASR Req $\Delta$ (%)
+MetaMath 7B 1,389 10.7 8.0 72.2
+WizardMath 13B 1,745 8.8 6.0 65.1
+MetaMath 70B 672 4.3 8.0 86.6
+GPT-3.5 456 2.8 10.0 90.9
+Target: GPT-4 5,000 29.3 10 -
+Table 5: Efficient Attack: A targeted approach to attack high-cost models.
+We leverage adversarial examples from cheaper models to attack GPT-4, achieving the same ASR while reducing up to 90% request calls.
 Efficient Attacks
-Scaling up the number of attacks results in a steady performance drop for all models.
-However, this scaling might not be viable for API-based models such as GPT-4 due to request limits and high costs.
-To circumvent this, we propose an efficient attack method for API-based models by using adversarial examples in a targeted manner.
+Scaling up the number of attacks leads to a consistent performance drop for all models.
+However, this may not be feasible for API-based models like GPT-4 due to request limits and high costs.
+To address this, we propose an efficient attack method for API-based models by using adversarial examples in a targeted manner.
 The idea is simple: we attack model $A$ (target model, GPT-4 in our case), with adversarial examples from a cheaper model $B$ (e.g., open-source, lower API cost).
@@
 We compare the results in Table 5 and observe a significant request reduction while achieving similar ASR.
-Models Request Call Cost (\$) ASR $\Delta$ Request (%)
-MetaMath 7B 1,389 10.7 8.0 72.2
-WizardMath 13B 1,745 8.8 6.0 65.1
-MetaMath 70B 672 4.3 8.0 86.6
-GPT-3.5 456 2.8 10.0 90.9
-Target: GPT-4 5,000 29.3 10 -
-Table 4: Efficient Attack A targeted approach to attack high-cost models.
-We leverage adversarial examples from cheaper models to attack GPT-4, achieving the same ASR while reducing up to 90% request calls.
-Figure 2: Human Evaluation (Left) We report the average score among three annotators on the adversarial examples from GPT-3.5.
-$M3$ has the highest score across all metrics, indicating our best generation method correctly generates contextually coherent problems that preserve original difficulty.
-Transferability (Right) We present the adversarial example transferability among all models by comparing each model against all other models.
-Compared to the math-tuned and production models, the weaker models such as Mistral 7B and LLama2 13B exhibit significant vulnerability and a strong correlation among them.
 5 Analysis and Discussion
-In this section, we conduct analyses and discussions to better understand our generation methods and the mathematical capabilities of LLMs.
-Human Evaluation
-We conduct a human evaluation to ensure the our generation are indeed valid.
-Three human evaluators are asked to assess (i) correctness: whether the answer to the modified problem is correct; (ii) coherence: whether the problem is contextually coherent; and (iii) similarity: whether the newly generated problem’s difficulty level is similar to the original one.
-We randomly select one adversarial example from 20 problems that GPT-3.5 failed on for GSM8K.
-The annotators made binary decisions to determine whether the criteria were met.
-We present the average of their scores on the left of Figure 2.
-Overall, all human evaluators agree on the correctness.
-$M3$ scores highly across all metrics, indicating that it preserves the original problem’s coherence and difficulty level.
-Coherence and similarity are relatively low for $M2$ and $M3$, suggesting that while they might perturb the model, such generation may not be useful for the educational context.
+In this section, we conduct analysis to better understand our generation methods and the mathematical capabilities of LLMs.
+Validation Through Human Evaluation
+We conduct a human evaluation to ensure the validity of our generated problems.
+We randomly select 30 problems that GPT-3.5 failed on for GSM8K and select 2 adversarial examples from each problem.
+Three evaluators are asked to assess (i) correctness: whether the answer to the modified problem is correct; (ii) coherence: whether the problem is contextually coherent; and (iii) similarity: whether the newly generated problem’s difficulty level matches the original.
+Evaluators make binary decisions to determine if the criteria are met, and Figure 2 shows the average scores.
+Overall, all evaluators agree on correctness.
+$M3$ scores highly across all metrics, indicating it preserves the original problem’s coherence and difficulty.
+Coherence and similarity scores for $M2$ and $M3$ are relatively low, suggesting their perturbations might not be useful in an educational context.
+Figure 2: Human Evaluation: The average score from three annotators.
+$M3$ achieves the highest scores across all metrics, indicating our best generation method correctly generates contextually coherent problems that preserve original difficulty.
 Transferability
-We investigate the transferability of adversarial examples among the models.
-We attack model $A$ with the adversarial examples from model $B$ and calculate the number of common adversarial examples between these two models.
-We present $M3$ result in the left of Figure 2.
-Interestingly, we observe that weaker models exhibit a high percentage of transferability, indicating a strong vulnerability correlation among these models.
-On the other hand, comparisons with strong models such as MetaMath 70B and GPT-3.5, tend to have a lower transferability rate.
+Following the previous work Zou et al.
+(2023); Zhou et al.
+(2023), we investigate the transferability of adversarial examples among the models.
+Specifically, We attack model $A$ with the adversarial examples from model $B$ and calculate the number of common adversarial examples between these two models.
+We present $M3$ result in the Figure 3.
+We observe that weaker models exhibit a high percentage of transferability, indicating a strong vulnerability correlation among them.
+On the other hand, strong models such as MetaMath 70B and GPT-3.5 tend to have a lower transferability rate.
 While it might suggest a form of resistance to adversarial examples that affect other models, it is not a measure of robustness as it could be that those models fail on entirely different examples.
@@
 To further understand these phenomena, we conduct regression analysis as shown below.
-Question Feature Analysis
-To further understand the limitations of LLMs, we conduct a regression analysis to investigate the effects of various features of MWPs on the correctness of the model’s predictions.
-We construct a total of 51 input features from 20,000 $M3$ generated problems, including features such as operation counts, value ranges, and node counts in the AST.
-The features are plotted against the correctness of the model, with correct being 1 and wrong being 0.
-We present the coefficients of the all LLMs with some selected features in Table 7 in the Appendix.
-We observe that the most positively and negatively correlated features vary across different models, suggesting that our adversarial examples exploit different weaknesses of each model.
-For example, models exhibit varying performance with different answer value ranges.
+Figure 3: Transferability: We present the adversarial example transferability (%) among all models by comparing each model against all other models.
+Compared to the math-tuned and production models, the weaker models such as LLama2 13B exhibit significant vulnerability and a strong correlation among them.
+Regression Feature Analysis
+To gain deeper insights into the limitations of LLMs on MWPs, we conduct a regression analysis investigating the relationships between various features of the problems and the correctness of the models’ predictions.
+We construct a set of 51 input features from 20,000 $M3$ generated problems, including features such as operation counts, answer value ranges, and node counts in the problem’s AST.
+By examining the coefficients of these features in predicting model correctness (see Table 8 and Figure 8 in the Appendix), we uncover interesting patterns that shed light on the limitations of LLMs:
+  • Varying Vulnerabilities Across Models We find that the most positively and negatively correlated features vary considerably across models, suggesting that our adversarial examples exploit distinct vulnerabilities.
+For instance, models exhibit divergent performance on problems with different answer value ranges.
 Mistral 7B and WizardMath 13B perform relatively well on problems with smaller answer values (e.g., in the range of [2, 8) and [8, 32)), while MetaMath 7B and Vicuna 13B show better performance on problems with answer values in the range of [32, 128).
-This suggests that different models may have learned to specialize in problems with specific range, possibly due to the varying distributions of the pre-training data.
-Overall, problems involving division and more operations tend to be more challenging for most LLMs.
-This confirms the intuition that complex problems requiring more mathematical operations are generally more difficult for LLMs to solve (Imani et al., 2023b).
-Visualizations are available in Figure 7 in the Appendix.
-Python Code Analysis
-While the majority of generated Python code adheres to correct problem-solving logic and produces answers correctly, to ensure that the generated code is correctly being converted into ASTs, we manually validate the generated ASTs and identify the following four types of possible errors: (1) Incorrect answers or code; (2) Use of loops, conditional, and comparison statements, and functions; (3) Use of expressions such as x = y // z + (y % z > 0); (4) Number misalignment, where the same number appears multiple times in the code or the problem.
-Encouragingly, for GSM8K and MultiArith, 96.5% and 92.5% of ASTs, respectively, are found to be valid and free from these errors.
-We discuss the methods of identifying and handling such errors in Appendix I.
-Why LLMs suffer From Such Simple Attack?
-While LLMs can be trained on vast amounts of math data, they may not have an inherent grasp of the underlying mathematical concepts and reasoning steps (Saxton et al., 2019; Lample & Charton, 2019).
+This observation hints that different models may have learned to specialize in problems with specific ranges, possibly due to variations in the distributions of their pre-training data.
+  • Complexity and Operation Types We observe that problems involving division and a higher number of operations tend to be more challenging for most LLMs (see Figure 8 in the Appendix for visualizations).
+For instance, problems requiring multiple division operations or a combination of different operations (e.g., addition, subtraction, multiplication, and division) are more likely to result in incorrect predictions.
+This aligns with the intuition that complex problems requiring more mathematical operations are generally more difficult for LLMs to solve (Imani et al., 2023b).
+  • Tokenization Choices and Numerical Reasoning: Recent research has also highlighted the impact of tokenization choices on LLMs’ numerical reasoning capabilities (Singh and Strouse, 2024).
+Our experiments reveal similar patterns, with LLama-based models tokenizing each digit individually, while GPT-3.5 and GPT-4 encode every three digits into a single token.
+This difference in tokenization may contribute to the observed consistent performance of GPT-3.5 to the number of tokens in the answer compared to the LLama-based models.
+Why LLMs suffer From Such Simple Attacks?
+While LLMs can be trained on vast amounts of math data, they may not have an inherent grasp of the underlying mathematical concepts and reasoning steps (Saxton et al., 2019; Lample and Charton, 2019).
 Instead, they likely rely on pattern and statistical associations learned from the training data to make predictions (Bender et al., 2021), which is also know as memorization (Carlini et al., 2022).
 Modifying the numbers in a problem may disrupt these learned patterns, causing the models to make errors.
-This is evidenced by the significant performance drops observed in models like WizardMath (89% to 20%) and Vicuna (76% to 4%) when presented with adversarial examples.
+This is evidenced by the significant performance drops observed in models like WizardMath (89% to 20%) and Vicuna (76% to 4%) when presented with adversarial examples (Table 1).
 Furthermore, test data contamination could also play a role - if the training corpus contains the test data, the models may be able to answer the original questions correctly but struggle to generalize to even subtle variations (Balloccu et al., 2024).
-6 Conclusion and Future work
-We introduce a novel approach to generate MWPs that challenge the LLM’s math-solving abilities, thereby addressing the urgent need for methods that ensure fair educational evaluations.
-By utilizing AST to create adversarial examples, we preserve problems’ original difficulty and coherence while significantly degrading both open- and closed-source LLMs and surpassing the previous rephrasing attack by an average of 60% ASR.
-We conduct a human evaluation to validate the adversarial examples and investigate universal attacks and attack transferability, proposing a cost-effective method to attack high-cost API-based models, achieving up to a 90% request reduction while maintaining the same performance.
-Our automatic regression analyses demonstrate that our adversarial examples exploit different weaknesses of each model, shedding light on the specific mathematical aspects that most influence individual LLM performance.
-This work not only paves the way for further investigation into the mathematical limitations among LLMs but also contributes to developing more robust and ethically sound educational tools to ensure fair evaluation, contributing to the ongoing discourse on the ethical use of LLMs in education.
-In this work, the correlation between the complexity that challenges LLMs and the actual difficulty perceived by humans has not been empirically validated.
-In future work, we aim to explore this and expand the methods to handle more complex problems.
+Generated Python Code and ASTs Analysis
+To ensure that the generated Python code is valid and can be correctly converted into ASTs, we manually examined 200 Python code and ASTs pairs from GSM8K and MultiArith, respectively.
+We identify the four types of errors:
+  • Incorrect answers or code: The generated code produces incorrect answers or contains logical errors that deviate from the problem statement.
+  • Use of unsupported constructs: The code includes loops, conditional and comparison statements, or user-defined functions, which are not supported by our AST conversion process.
+  • Complex expressions: The code contains expressions that are difficult to convert into ASTs, such as x = y // z + (y % z > 0).
+  • Number misalignment: The same number appears multiple times in the code or the problem, leading to inconsistencies between the generated code and the problem statement.
+Encouragingly, we find that 96.5% and 92.5% of the generated ASTs for GSM8K and MultiArith, respectively, are valid and free from these errors.
+We discuss this process in more details in the Appendix H.
+The high percentage of valid code and ASTs pair indicates that our code generation approach is indeed reliable.
+6 Conclusion
+We introduce a novel method to generate adversarial MWPs using ASTs.
+Our approach effectively challenges the mathematical problem-solving abilities of LLMs while maintaining the original difficulty and coherence of the problems.
+The generated adversarial examples significantly degrade the performance of both open- and closed-source LLMs, surpassing previous attacks by an average of 62% ASR.
+We validate our adversarial examples through human evaluation and investigate universal attacks and transferability, proposing a cost-effective method to attack high-cost API-based models with up to a 90% reduction in requests.
+Automatic regression analyses reveal distinct weaknesses of different models when solving MWPs.
+Our work contributes to the development of fair and robust educational tools, ensuring ethically sound evaluations and promoting the responsible use of LLMs in education.
+7 Limitations
+This work has not empirically validated the correlation between the complexity of generated problems and the actual difficulty perceived by human students.
+Further investigation is needed to establish this correlation.
+As a proof of concept, our method demonstrates the feasibility of generating meaningful and useful adversarial examples to LLMs, at least at the difficulty levels represented by the GSM8K and MultiArith datasets.
+However, it may not generalize well to math problems with significantly different types or difficulty levels.
+For example, very simple arithmetic problems (e.g., 3 + 2 = ?) might never have a successful attack on LLMs specifically trained for math, while advanced problems requiring proofs or concepts without numeric manipulation (e.g., prove newton’s binomial theorem) may not be suitable for our code generation procedure.
+As LLMs continue to improve, it is possible that they may become more robust, which is a challenge faced by all adversarial attack methods in general.
+We plan to address these limitations in future work by exploring a wider range of math problems and investigating other attack vectors as LLMs evolve.
 8 Ethics Statement
@@
 This approach risks amplifying the disparities between institutions or individuals with differing levels of technological access.
+9 Acknowledgements
+We are very grateful to Sandy Yeung and Jack Goffinet for the helpful initial project discussion.
+Thank you to the anonymous reviewers for their feedback.
+The research reported here was supported by the Learning Engineering Virtual Institute funded by leading education philanthropists and organizations through Grant G-23-2137070, to the University of Florida and its partner institutions.
+The opinions expressed are those of the authors and do not represent the views of the University of Florida, the partner institutions, or those of the philanthropists and organizations.
 Appendix A Prompt Templates
 A.1 Code Generation Prompt
-Code Generation Prompt Write a Python script that contains a step-by-step solution for a given math problem and its answer.
+Write a Python script that contains a step-by-step solution for a given math problem and its answer.
 Start the script by defining variables corresponding to the quantities mentioned in the problem.
@@
 A.2 Zero-shot CoT
-Evluation Prompt Solve a math problem. The solution ends with “the answer is (a number)” like “the answer is 1”. Question: {problem} Answer: Let’s think step by step.
-Appendix B Rephrasing Example
+Solve a math problem. The solution ends with “the answer is (a number)" like “the answer is 1". Question: {problem} Answer: Let’s think step by step.
+Appendix B A Rephrasing Attack Example
 Figure 4: A rephrasing attack example from Zhou et al.
@@
 We associate the number in the original question with the variable node so that a change in the variable node will also change the number in the question.
-If a number appears multiple times in the code and the problem, then we associate them based on the specific procedures described in Section I.4.
+If a number appears multiple times in the code and the problem, then we associate them based on the specific procedures described in Section H.4.
 Appendix D Generation Details
@@
 MetaMath 7B 91.1 79.3 13.0
+Llama 3 8B 22.0 30.0 0.0
 Llama-2 13B 2.3 8.3 0.0
@@
 GPT-3.5 91.1 75.7 16.9
-Table 6: Rephrasing Attack: As a baseline, we present the performance of RobustMath Zhou et al.
-(2023).
-While rephrasing the problem results in lower accuracy for 4 out of 7 models, a few models show improved accuracy with the rephrased version.
+Table 7: While rephrasing the problem results in lower accuracy for 4 out of 7 models, a few models show improved accuracy with the rephrased version.
 This suggests that rephrasing the MWPs may not generalize effectively across different LLMs.
 Appendix F Incorrect Variant Count Distribution
-Figure 4: We present the distribution of the number of incorrect adversarial examples in different buckets.
-For strong models (e.g., GPT-3.5-Turbo, MetaMath 70B), around half of the problems have zero incorrect adversarial examples, while for weak models (e.g., Llama-2 13B, CodeLlama 34B), most problems has more than 90 incorrect adversarial examples.
+Figure 5: We present the distribution of incorrect adversarial examples in different buckets.
+For strong models (e.g., GPT-3.5-Turbo, MetaMath 70B), around half of the problems have zero incorrect adversarial examples, while for weaker models (e.g., Llama-2 13B, CodeLlama 34B), most problems have more than 90 incorrect adversarial examples.
 Appendix G Number of Generation
-Figure 5: Possible Generations: with 30,000 attempts for each given constraint, we calculate the average number of adversarial examples generated for each problem.
-Appendix H Number of Attacks
-Figure 6: We report the model performance given various number of attacks. For several models, only 10 attacks are enough to degrade the performance.
-Appendix I Code Analysis
-I.1 Incorrect Answer from the Dataset
-There are instances where the dataset-provided answers might be wrong.
-An example of such an instance is the following question from the MultiArith dataset: ”Paige’s team won their dodgeball game and scored 41 points total.
-If Paige scored 11 of the points and everyone else scored 6 points each, how many players were on her team?”, which the dataset provided 5 as the answer.
+Figure 6: With 30,000 attempts for each given constraint, we calculate the average number of adversarial examples generated for each problem.
+Appendix H Code Analysis
+H.1 Incorrect Answer from the Dataset
+Given the execution steps in code are almost always deterministic, we are able to identify this error effectively.
+There are also instances where the dataset-provided answers might be wrong.
+An example of such an instance is the following question from the MultiArith dataset: "Paige’s team won their dodgeball game and scored 41 points total.
+If Paige scored 11 of the points and everyone else scored 6 points each, how many players were on her team?", which the dataset provided 5 as the answer.
 However, the correct answer should be 6, as Paige is on the team as well.
@@
 To automatically detect such instances, the value of the final answer node will be compared to the answer given by the dataset, and the question will be filtered out for further generations of problems if the two answers don’t match.
-I.2 Incorrect Answer from GPT-4
+In total, we found only 1% and 2.5% generated codes have incorrect logic or answers in GSM8K and MultiArith, respectively.
+H.2 Incorrect Answer from GPT-4
 Similarly, there are also instances where GPT-4 generated codes that provided the wrong answer to the given question.
-One such question is from the GSM8K dataset: ”Jen decides to travel to 3 different countries.
+One such question is from the GSM8K dataset: "Jen decides to travel to 3 different countries.
 He has to pay \$400 for the supplies he needs, in total.
 The tickets for travel cost, in total, 50% more than the supplies.
-How much does travel cost?” The correct answer should be $400+400\times 1.5=1000$, but the code generated by GPT-4 outputted \$600.
+How much does travel cost?" The correct answer should be $400+400\times 1.5=1000$, but the code generated by GPT-4 outputted \$600.
 Similar to the previous error, instances where the code outputs the wrong answers will be detected by comparing them to answers given by the dataset, and the instances will be discarded from generating math problems.
-I.3 Contains Control Flow Statements
-Although specified in the prompt, GPT-4 might generate codes that include control flow statements (including ”if”, ”for”, and ”while” statements).
+H.3 Contains Control Flow Statements
+Although specified in the prompt, GPT-4 might generate codes that include control flow statements (including "if", "for", and "while" statements).
 Codes that contain control flow statements are considered errors and discarded because the codes are usually only applicable to the original instance of the question.
 Furthermore, the code might not halt for some combinations of numbers, resulting in no answers associated with the generated instances.
-I.4 Number Misalignment
+H.4 Number Misalignment
 To ensure the correct generation of new questions, numbers in the original question need to be associated with numbers in the code.
@@
 Then, random numbers are generated to combine the equivalences of correspondences.
-For example, given that the question is ”Mary has 5 apples and 5 oranges, how many fruits does Mary have?”.
-Although the two ”5”s correspond to different representations in the question, due to the communicative nature of addition, the two ”5”s are interchangeable in the syntax tree.
+For example, given that the question is "Mary has 5 apples and 5 oranges, how many fruits does Mary have?".
+Although the two "5"s correspond to different representations in the question, due to the communicative nature of addition, the two "5"s are interchangeable in the syntax tree.
 Thus, the two pairs of correspondence for this question can be combined, and there is no ambiguity in the grounding.
@@
 The correspondence that has the same answer will be used, and if no correspondence has the correct answer, the question will be discarded.
-An example of this type of question is: ”The pizzeria sells small pizzas for \$2 and large pizzas for \$8.
+An example of this type of question is: "The pizzeria sells small pizzas for \$2 and large pizzas for \$8.
 They sold \$40 in pizzas.
-If they sold 8 small pizzas, how many large pizzas did they sell?”.
+If they sold 8 small pizzas, how many large pizzas did they sell?".
 One valid and distinguishing sequence of numbers for the question is $(2,3,31,5)$, which will result in different answers $5$ and $7$.
@@
     There is a number $c_{i}$, such that $C_{q}(c_{i})>C_{c}(c_{i})$. Similar to the reasoning for the first case, these kinds of questions are eliminated for problem generations.
+Appendix I Number of Attacks
+Figure 7: We report the model performance given various number of attacks. For several models, only 10 attacks are enough to degrade the performance.
 Appendix J Feature Analysis